### v1

# ReasoningRank: Teaching Student Models to Rank through Reasoning-Based Knowledge Distillation

Abstract
Reranking documents based on their relevance to a given query is a critical task in information retrieval. Traditional reranking methods often focus on improving the initial rankings but tend to lack transparency, failing to explain why one document is ranked higher than another. In this paper, we introduce ReasoningRank, a novel reranking approach that enhances clarity by generating two types of reasoning: explicit reasoning, which explains how a document addresses the query, and comparison reasoning, which justifies the relevance of one document over another. We leverage large language models (LLMs) as teacher models to generate these explanations and distill this knowledge into smaller, more resource-efficient student models. While the student models may not outperform LLMs in speed, they significantly reduce the computational burden by requiring fewer resources, making them more suitable for large-scale or resource-constrained settings. These student models are trained to both generate meaningful reasoning and rerank documents, achieving competitive performance across multiple datasets, including MSMARCO and BRIGHT. Experiments demonstrate that ReasoningRank improves reranking accuracy and provides valuable insights into the decision-making process, offering a structured and interpretable solution for reranking tasks.  [^1] Our code and data are available here https://anonymous.4open.science/r/Distillation-Is-All-You-Need-C98D/README.md
ReasoningRank: Teaching Student Models to Rank through Reasoning-Based Knowledge Distillation
Yuelyu Ji1, Zhuochun Li1, Rui Meng2, Daqing He1 1School of Computing and Information, University of Pittsburgh, 2Salesforce Research {yuj49,zhl163, dah44}@pitt.edu, ruimeng@salesforce.com
1 Introduction
The ability to effectively rank and rerank documents is critical in information retrieval (IR), where the quality of search results significantly impacts user satisfaction Pradeep et al. (2023b); Sun et al. (2023). Traditional reranking approaches often improve initial rankings provided by retrieval systems but pay limited attention to the reasoning behind ranking decisions. Recent studies have demonstrated that incorporating reasoning into the reranking process can enhance performance and interpretability Pradeep et al. (2023b); Liu et al. (2024); Yu et al. (2024c); Sun et al. (2023). By offering explanations that clarify why certain documents are prioritized over others, these methods help users understand and trust the search results.
However, most existing reranking models do not distinguish between two key types of reasoning: explicit reasoning, which explains how a document matches a query, and comparative reasoning, which justifies why one document ranks higher than another. This limitation can reduce ranking effectiveness, particularly in contexts where a subtlety understanding of document relevance is crucial.
Figure 1: Illustration of the two types of reasoning: explicit reasoning provides direct answers to the query, while comparison reasoning evaluates the relative relevance between documents. The LLM generates these explanations to enhance the interpretability of the ranking process.
In this paper, we propose ReasoningRank, a reranking framework that leverages both explicit and comparative reasoning to provide more detailed and transparent document rankings. Our approach is based on knowledge distillation, where a large, powerful teacher model (an LLM) generates high-quality explanations and rankings. These explanations are then distilled into smaller, more computationally efficient student model that perform the reranking tasks. Our contributions are threefold:
  • We introduce a reranking approach that separates explicit and comparative reasoning, enhancing clarity of document rankings.
  • We demonstrate that student models can learn from LLM-generated data, achieving high reranking performance while remaining computationally efficient.
  • We provide an open-source implementation to facilitate further research and applications in information retrieval.
The remainder of this paper is organized as follows. Section 2 reviews related work. Section 3 presents the proposed ReasoningRank framework. Section 4 not only presents the quantitative performance metrics but also provides a thorough discussion on the implications of the results, analyzing how the reasoning strategies impact the ranking performance and offering insights into the model’s interpretability. Finally, Section 6 concludes the paper and discusses future work.
2 Related Work
The evolution of reranking techniques has significantly advanced with the rise of neural networks, particularly transformers Vaswani (2017); Yu et al. (2024b, a); Zheng et al. (2024); Zhang et al. (2024); Shen et al. (2024); Fan and Tao (2024); Fan et al. (2024). Early models like monoBERT Nogueira et al. (2019) utilized BERT Devlin (2018) in a cross-encoder setup, effectively capturing contextual dependencies between queries and documents. Further models such as monoT5 Nogueira et al. (2020) and InPars Bonifacio et al. (2022) continued to push reranking performance boundaries, especially on datasets like MS MARCO Bajaj et al. (2016). However, these models rely heavily on large-scale fine-tuning and task-specific data, limiting their scalability and adaptability to new domains or tasks.
Figure 2: Overview of explicit and comparison reasoning prompts for document ranking. Explicit reasoning explains how a document matches a query, while comparison reasoning evaluates the relative relevance between documents.
Figure 3: Overview of the ReasoningRank framework. The teacher model generates explicit and comparative reasoning, which is used to train student models capable of reproducing the reranked order and generating explanations.
Early ranking systems primarily focused on pointwise methods, where each document’s relevance was judged in isolation Liu et al. (2009). However, pairwise approaches like duoBERT Pradeep et al. (2021) were introduced to compare pairs of documents, determining which document was more relevant. While this improved the relative ranking of documents, it still missed the larger picture of the entire ranked list. To address this, listwise models such as RankGPT Sun et al. (2023) and its distilled versions like RankVicuna Pradeep et al. (2023a) considered the entire ranked list during training, allowing for more holistic improvements in ranking accuracy. However, even these methods have limitations when it comes to explaining why a particular document ranks higher than another.
The Learning to Rank framework introduced several foundational approaches in information retrieval, including pointwise, pairwise, and listwise methods, which form the basis of modern neural ranking models Liu et al. (2009). Pointwise approaches focus on predicting the relevance score of individual documents in isolation, while pairwise methods aim to indicate the relative ranking between document pairs. On the other hand, listwise approaches focus on directly optimizing the entire ranking list, offering better overall accuracy by considering the global structure of the ranking. Despite these advances, most early methods lacked clarity that modern applications demand.
Recent work has also highlighted the importance of natural language explanations in improving clarity and trust in reranking decisions. Models such as Explain-then-Rank Zhuang et al. (2023) have focused on generating explanations that describe how a document matches a query. However, these approaches often only address explicit reasoning (i.e., document-query matching) and do not account for comparative rationale (i.e., why one document ranks higher). Our work builds on these efforts by incorporating both types of reasoning, enhancing the overall clarity and coherence of the explanations. This dual focus provides more precise insights into the ranking process and improves the model’s ability to deliver accurate and logically consistent rankings Doshi-Velez and Kim (2017); Lipton (2018); Mo et al. (2024); Yu et al. (2024b, a).
Knowledge distillation has become an essential technique in training smaller, more efficient student models from larger teacher models Hinton et al. (2015) to further reduce the computational demands of large reranking models. Knowledge distillation has been particularly effective in reranking tasks, as demonstrated by models like RankVicuna Pradeep et al. (2023a), which distilled the reranking capabilities of RankGPT into a smaller, more efficient model without significant loss of performance. Our approach builds on this by applying distillation to both the reasoning and ranking components, ensuring that our student models not only maintain high performance but also offer explainability in their reranking decisions Qin et al. (2023); Dai et al. (2022); Xu et al. (2024b, a).
By integrating these advancements, our ReasoningRank framework improves upon existing methods by offering a dual reasoning strategy—explicit and comparative reasoning—within a computationally efficient model. Our approach not only enhances ranking accuracy but also provides clear and easy-to-understand explanations for the ranking decisions, addressing fundamental limitations in performance seen in prior work.
3 Method
The overview of the architecture of the ReasoningRank model is shown in Figure 3. We propose ReasoningRank, a framework that integrates explicit and comparative reasoning within a unified reranking approach.
3.1 Teacher Model with Dual Reasoning
Given a query $q$ and a set of candidate documents $\{d_{1},d_{2},\dots,d_{n}\}$, the teacher model processes these inputs to generate two types of reasoning and rerank order. The teacher model can generate high-quality reasoning that combines both explicit and comparative aspects, and the examples can be seen in Figure 2: Explicit Reasoning: The model explains how each document directly addresses the query, focusing on the relevance and specificity of the content. Comparative Reasoning: The model assesses the relative relevance between documents, explaining why one document should be ranked higher based on content.
3.2 Student Model with Knowledge Distillation
Deploying large language models (LLMs) for reranking is often impractical due to their high computational cost. To address this, we designed the student model to replicate the teacher’s performance while being more efficient for real-world deployment. The innovation lies in using knowledge distillation: the student learns from the teacher’s detailed reasoning and ranking outputs, retaining accuracy and interpretability with significantly lower computational demands. This approach ensures the student can perform complex reranking tasks efficiently, making it ideal for large-scale use.
Architecture and Training:
Given a query $q$ and a set of candidate documents $\{d_{1},d_{2},\dots,d_{n}\}$, the model processes these inputs to generate both reasoning outputs and ranking scores.
The model generates two outputs: Reasoning Outputs: The model generates textual reasons $\mathbf{r}_{i}$ explaining the relevance of each document $d_{i}$ to the query $q$. Ranking Scores: The model simultaneously generates ranking scores $\mathbf{s}_{i}$ for each document, indicating its relevance to the query.
3.2.1 Training Objective and Loss Functions
To optimize the performance of the student model, we define three distinct loss functions, each corresponding to different aspects of the model. First, using pairwise loss to optimize the relative ranking between pairs of documents, ensuring that the more relevant document is ranked higher. The pairwise loss is defined as:
$\mathcal{L}_{\text{pairwise}}=\sum_{(i,j)\in P}\max(0,1-(\mathbf{s}_{i}-\mathbf{s}_{j}))$
where $\mathbf{s}_{i}$ and $\mathbf{s}_{j}$ represent the predicted relevance scores of documents $d_{i}$ and $d_{j}$, and $P$ is the set of document pairs with known relevance relations.
Next, the listwise loss optimizes the overall ranking of the document list. It uses Kullback-Leibler (KL) divergence to measure the difference between the predicted ranking distribution and the target ranking distribution:
$\mathcal{L}_{\text{listwise}}=D_{\text{KL}}(Q(\mathbf{z})\parallel P(\mathbf{s}))$
Where $P(\mathbf{s})=\text{softmax}(\mathbf{s}_{i})$ is the predicted ranking distribution based on the model’s output, and $Q(\mathbf{z})=\text{softmax}(\mathbf{z}_{i})$ is the target ranking distribution based on true labels.
Finally, the generation loss optimizes the quality of the textual reasoning generated by the model. Based on the cross-entropy loss between the generated reasoning and the ground-truth reasoning:
$\mathcal{L}_{\text{generation}}=-\sum_{i=1}^{n}\mathbf{y}_{i}^{\text{gen}}\log(\text{softmax}(\mathbf{r}_{i}))$
where $\mathbf{r}_{i}$ represents the generated reasoning logits for document $d_{i}$, and $\mathbf{y}_{i}^{\text{gen}}$ is the ground truth reasoning label.
3.3 Final Training Objective
To achieve balanced training, we introduce learnable weights $\alpha$, $\beta$, and $\gamma$ to adjust the contributions of pairwise, listwise, and generation losses, respectively. The final loss function is:
$\mathcal{L}=\alpha\times\mathcal{L}_{\text{pairwise}}+\beta\times\mathcal{L}_{\text{listwise}}+\gamma\times\mathcal{L}_{\text{generation}}$
Where $\alpha+\beta+\gamma=1$, ensuring a balanced combination of the three loss functions during training. With this design, the student model efficiently performs ranking tasks and provides transparent reasoning for its decisions, enhancing both transparency and performance.
4 Experiments and Results
In this section, we evaluate the performance of our proposed ReasoningRank framework across various information retrieval tasks. We introduce the datasets and evaluation metrics used in our experiments (§4.1). We then detail the experimental setup and baseline methods for comparison (§4.3). Finally, we present a comprehensive analysis of the experimental results, including comparisons with existing methods in (§4.4). We performed paired t-tests across all queries in each dataset to assess the statistical significance of the improvements. The annotations in Tables 1 and 2 indicate the significance levels.
4.1 Datasets and Evaluation Metrics
In addition to MSMARCO Bajaj et al. (2016) and BEIRThakur et al. (2021), we utilize the BRIGHT Su et al. (2024) to further evaluate our model’s ability to handle reasoning-intensive retrieval tasks.
Evaluation Metrics
We use the following evaluation metrics to measure the performance of our models:
Normalized Discounted Cumulative Gain (NDCG@k): Measures the ranking quality of the retrieved documents, with higher emphasis on top-ranked documents. We report NDCG at rank positions 5 and 10.
BLEU, ROUGE-L, and Coherence Score: Used for evaluating the quality of generated reasons in reasoning tasks, assessing both lexical overlap and logical consistency.
4.2 Teacher Models
We utilize large language models (LLMs) as teacher models to generate reasoning annotations: GPT-4: A state-of-the-art LLM known for its strong reasoning and language understanding capabilities. ClaudeThe : An advanced LLM developed by Anthropic, offering competitive performance in reasoning tasks, we used the version claude-3-5-sonnet-20240620. GeminiTeam et al. (2023): A commercial LLM used for comparison purposes.We used the version gemini-1.5-flash.
4.2.1 Overall Performance on Benchmark Datasets
Table 1 summarizes the NDCG@5 performance of different models GPT-4 consistently achieves the best results, with 77.69% on DL19 and 85.27% on TREC-COVID when using both comparison and explicit reasoning prompts, showcasing its robust handling of complex queries.
Adding different reasoning prompts to the LLMs enhances their performance across various datasets. For instance, GPT-4 with comparison and explicit reasoning outperforms its baseline (without reasoning) on DL19 by 2.10%. Similar trends are observed with Claude and Gemini, indicating that incorporating reasoning prompts in different LLMs is beneficial.
To confirm the statistical significance of the improvements, we conducted paired t-tests comparing models with reasoning prompts to their respective baselines. The results, indicated by and in Table 1, show that the enhancements are significant, with p-values less than 0.05.
Given GPT-4’s overall solid performance and the boost from reasoning prompts, we select it as the teacher model to distill its enhanced reasoning abilities into our student models.
Models DL19 DL20 Covid Touche News NFCorpus Robust04 DBPedia
BM25 50.58 47.96 59.47 44.22 39.52 30.75 40.73 31.83
Rank GPT-4 Models
Rank GPT-4 (without reasoning) 75.59 70.56 85.51 38.47 57.55 36.62 57.55 47.12
+ Explicit reasoning 75.32 69.93 83.27 38.94 55.87 36.08 52.89 46.98
+ Comparison reasoning 76.25 71.07 84.62 38.23 56.49 36.52 56.53 48.15
+ Comparison & Explicit 77.69 73.24 85.27 38.29 58.43 36.33 58.63 49.52
Claude Models
Claude (without reasoning) 71.96 68.81 82.77 36.12 54.35 36.36 50.78 42.51
+ Explicit reasoning 70.59 68.23 82.22 35.92 55.43 35.59 52.38 42.51
+ Comparison reasoning 71.34 69.75 83.31 36.43 55.11 36.71 51.53 43.23
+ Comparison & Explicit 72.13 69.95 84.03 37.33 55.52 36.36 52.67 44.91
Gemini Models
Gemini (without reasoning) 71.10 68.45 82.95 35.62 53.65 35.49 50.12 42.92
+ Explicit reasoning 69.08 67.02 81.55 34.83 52.87 36.92 51.31 43.92
+ Comparison reasoning 70.17 68.24 82.41 35.91 53.35 35.42 51.12 42.85
+ Comparison & Explicit 71.35 68.57 83.47 36.97 53.80 36.12 51.78 43.67
Table 1: NDCG@5 performance (in percentage) for different models with various reasoning prompts across multiple datasets. Bold numbers indicate the best performance within each group of models. indicates statistically significant improvement over the base model without reasoning prompts (p $<$ 0.05). indicates statistically significant improvement over both the base model and the single reasoning prompt variants (p $<$ 0.05).
Models DL19 DL20 Covid Touche News NFCorpus Robust04 DBPedia
Student Models
Basic w/o reasoning 73.23 70.80 79.94 35.10 51.90 34.60 52.82 46.31
With explicit reasoning 74.54 70.11 79.37 34.25 52.81 35.45 50.75 46.95
With comparison reasoning 74.12 72.25 80.10 34.00 53.10 36.63 52.11 47.21
Comparison & Explicit 75.39 72.36 84.59 36.22 53.81 36.37 53.54 47.85
Baseline Models
DeBERTa Sun et al. (2023) 68.53 64.18 73.42 32.13 50.24 33.67 49.23 45.39
MonoT5 Nogueira et al. (2020) 74.54 70.44 79.95 34.07 46.03 35.21
RankVicuna Pradeep et al. (2023a) 71.12 68.66 67.14 48.73 38.53 55.67 35.29
RankZephyr Pradeep et al. (2023b) 72.23 70.53 85.12 36.46 53.33 38.92 60.72 35.52
APEER Jin et al. (2024) 74.57 72.25 83.86 35.25 52.11 33.41 56.03 46.13
Table 2: NDCG@5 performance (in percentage) for student models and baseline comparisons across multiple datasets. indicates statistically significant improvement over the baseline DeBERTa model (p $<$ 0.05). Bold indicates the best performance among all models for each dataset. Entries marked with "–" indicate that results are not available for that dataset.
4.3 Student Models
Experimental Setup
We implement our models using LoRA (Low-Rank Adaptation) on 32GB V100 GPUs to fine-tune the student models on large datasets efficiently. Based on the LLaMA 3.1 8B architecture, the student models fine-tuned with different reasoning prompts derived from the teacher models.
4.3.1 Performance Comparison with Baseline
Our student models, trained with explicit and comparison reasoning, achieve competitive performance compared to the previous state-of-the-art student models. As shown in Table 2,the improvements over the baseline models are statistically significant, with p-values less than 0.05 (indicated by ).
BM25: A traditional term-based retrieval model using TF-IDF, known for its simplicity and efficiency in keyword-based retrieval.
RankGPT DeBERTa Sun et al. (2023): A transformer model using disentangled attention to better capture long-range dependencies, improving reranking by understanding query-document context.
MonoT5 Nogueira et al. (2020): A T5-based model treating reranking as a sequence generation task, leveraging its encoder-decoder architecture for better relevance scoring.
RankVicuna, RankZephyr and APEER Pradeep et al. (2023a, b); Jin et al. (2024): RankVicuna is a distilled version of RankGPT, designed for zero-shot listwise reranking with fewer parameters, balancing efficiency and performance.RankZephyr a state-of-the-art zero-shot listwise reranking model that adapts RankGPT prompts for robust performance with minimal task-specific training. APEER share the same structure with RankZephyr but have an automated prompt generation for large language models.
4.4 ReasoningRank Student Model vs. Baseliness
Our ReasoningRank student model demonstrates strong performance across various datasets, often outperforming the baseline models. For instance, on the DL19 dataset, our student model achieves an NDCG@5 of 75.39%, surpassing DeBERTa (68.53%), RankVicuna (71.12%), and RankZephyr (72.23%). Similarly, on the DL20 dataset, our model attains 72.36%, outperforming DeBERTa (64.18%) and RankVicuna (68.66%).
On some datasets, our model achieves the best performance among all models. For example, on the News dataset, our student model reaches an NDCG@5 of 53.81%, slightly outperforming RankZephyr (53.33%) and significantly better than DeBERTa (50.24%).
However, there are cases where other models perform better. On the Touche dataset, RankVicuna achieves an NDCG@5 of 48.73%, which is higher than our student model’s 36.22%. This may be due to RankVicuna’s design for zero-shot listwise reranking, which could be particularly effective for argument retrieval tasks represented in Touche. Similarly, on the Robust04 dataset, RankZephyr attains a higher NDCG@5 of 60.72% compared to our model’s 53.50%.
These variations suggest that while our model performs well, certain models may have advantages in specific domains or datasets due to their architectures or training strategies. For example, RankVicuna’s strong performance on Touche might be attributed to its effectiveness in handling argumentation and persuasive texts, which are prevalent in that dataset. RankZephyr’s higher score on Robust04 could be due to its ability to leverage zero-shot capabilities effectively in that particular dataset.
4.4.1 Performance on Reasoning-Intensive Tasks (BRIGHT Dataset)
The BRIGHT dataset is specifically designed to evaluate models on reasoning-intensive retrieval tasks, simulating real-world scenarios where complex reasoning is required to determine document relevance. Table 3 shows the NDCG@5 results on the BRIGHT dataset. In the Earth Science domain, the student model achieves an NDCG@5 of 27.0%, closely approaching the teacher model’s 31.0%. Similarly, in Psychology, the student model attains 26.0%, compared to the teacher’s 28.9% and BM25’s 11.6%.
Domain BM25 Teacher Student GPT-4 COT
Bio. 17.2 24.3 21.0 28.0
Earth. 24.2 31.0 27.0 37.1
Econ. 13.2 19.0 16.5 22.1
Psy. 11.6 28.9 26.0 34.2
Rob. 12.0 16.0 14.5 19.0
Stack 16.0 21.0 18.5 24.2
Sus. 12.4 18.0 16.2 21.9
Leet. 22.1 27.6 25.0 31.4
Pony 8.5 5.9 5.2 6.7
AoPS 6.0 7.6 6.8 9.2
TheoQ. 9.4 18.5 15.3 23.0
TheoT. 3.5 17.0 14.2 21.7
Average 13.1 19.6 17.5 23.2
Table 3: NDCG@5 results on the BRIGHT dataset.
4.4.2 Quality of Generated Reasons
We assess the quality of the generated reasons using BLEU and ROUGE-L as presented in Table 4 compared with the teacher model’s reason. The baseline model here represents the model trained without the rank loss function and has the same input as ours. On the DL19 dataset, the student model achieves a BLEU score of 21.4, compared to the baseline’s 18.0, and a ROUGE-L score of 36.8 versus 32.5. Similar improvements are observed on DL20, where the student model attains BLEU and ROUGE-L scores of 24.0 and 38.2, respectively, surpassing the baseline’s 19.5 and 34.0.
Dataset Model BLEU ROUGE-L
DL19 Baseline 18.0 32.5
Ours 21.4 36.8
DL20 Baseline 19.5 34.0
Ours 24.0 38.2
Covid Baseline 16.0 30.0
Ours 18.9 34.5
Touche Baseline 17.5 33.0
Ours 20.5 37.1
News Baseline 20.0 35.0
Ours 22.7 39.3
Table 4: Evaluation of reason generation using BLEU, and ROUGE-L scores for baseline and student models.
4.4.3 Ablation Study: Impact of Reasoning Strategies
Results
In Table 5, by removing comparison reasoning results in a performance drop on DL19 from 75.39 to 74.12 (NDCG@5). Similarly, removing explicit reasoning causes a performance decrease on DL20 from 72.36 to 72.25 (NDCG@5).
Analysis
Both reasoning strategies contribute to performance improvements. The ablation study shows that omitting either reasoning component results in a noticeable decline in performance, highlighting their complementary roles in improving ranking effectiveness.
Model DL19 NDCG@5/@10 DL20 NDCG@5/@10
Comparison & explicit 75.39 / 73.85 72.36 / 71.22
w/ explicit reasoning 74.54 / 73.18 70.11 / 70.97
w/ comparison reasoning 74.12 / 72.94 72.25 / 70.43
Basic w/o reasoning 73.23 / 72.12 70.82 / 69.83
Table 5: Ablation study: Impact of explicit and comparison reasoning on NDCG@5/@10 for DL19 and DL20.
4.4.4 Effects of Training Data Size Across Student Models
Figure 4 i shows the performance of small models trained on 100, 1,000, and 2,000 examples. Contrary to expectations, increasing data size from 1,000 to 2,000 does not consistently improve performance and sometimes slightly reduces it. We observed similar trends across various small models, including Mistral and LLaMA variants, highlighting that adding more data does not always guarantee better performance.
Figure 4: Student model performance for different training data sizes on DL19 and DL20.
5 Discussion
Our experiments demonstrate that the ReasoningRank framework effectively enhances document reranking by incorporating explicit and comparative reasoning. The student models achieve competitive results with significantly reduced computational resources compared to large teacher models. The statistically significant improvements confirm the effectiveness of reasoning strategies in enhancing model performance.
An important aspect of ReasoningRank is its ability to not only rank documents but also generate coherent and relevant reasoning for its rankings. This dual capability suggests that the model could be applied to other tasks that require both ranking and reasoning generation.
While our current work focuses on document reranking in information retrieval, the approach of integrating reasoning generation with ranking decisions could be extended to other domains. For example, in recommendation systems, the model could rank products or services while providing explanations for its recommendations, thereby improving user trust. Similarly, in question answering or summarization tasks, generating reasoning alongside answers or summaries could enhance understandabilityJiang et al. (2024).
However, we acknowledge that our paper does not evaluate the model’s generative abilities on tasks beyond document reranking. This is a limitation in assessing the full scope of the model’s applicability. Future work should explore how the model performs in different contexts to verify its generalizabilityMcElfresh et al. (2022); Rahmani et al. (2022).
6 Conclusion
In this paper, we introduced ReasoningRank, a novel approach to document reranking that incorporates explicit and comparative reasoning. Our method leverages the power of large language models to generate interpretable reasons, which are then distilled into smaller, more efficient student models. Our experiments demonstrate the effectiveness of this approach, achieving high performance across multiple datasets while maintaining interpretability.
Limitations
While our approach shows strong performance, it has several limitations. First, the reliance on large language models for generating reasoning limits the applicability of our method in resource-constrained environments. Second, the quality of the generated reasons heavily depends on the quality of the prompts, which requires careful crafting. Future research should address these limitations by exploring more efficient models and automated prompt-generation techniques.
Ethics Statement
This work enhances the transparency and accessibility of reasoning capabilities in AI by distilling knowledge from LLMs to smaller models. While this improves efficiency, we acknowledge the ethical concerns of bias in LLMs and the need for careful mitigation in real-world applications.
Appendix A Student Training Parameters
Hyperparameter Value
low_cpu_fsdp True Interpretability
run_validation True
batch_size_training 1
context_length 1024
gradient_accumulation_steps 1
gradient_clipping True
gradient_clipping_threshold 1.0
num_epochs 3
max_train_step 0
max_eval_step 0
learning rate 1e-4
weight_decay 0.0
seed 42
use_fp16 True
use_peft True
freeze_layers False
Table 6: Training configuration hyperparameters.
Appendix B Teacher Model Parameters
Hyperparameter Value
Model GPT
Temperature 1.0
Top p 0.9
Table 7: Teacher model hyperparameter settings.
Appendix C NDCG@5 for Different Models
Table 8: NDCG@5 for different models and training data sizes on various datasets.
Model Train Size DL 19 DL 20
Mistral v0.1 100 69.69 76.33
1000 71.03 77.06
2000 70.63 73.99
Mistral v0.3 100 69.26 79.55
1000 71.56 79.14
2000 69.86 76.96
LLaMA 3.1 8B 100 69.63 75.85
1000 72.87 77.03
2000 73.37 77.60
LLaMA 3 8B 100 69.57 77.15
1000 73.76 75.35
2000 71.75 75.55
LLaMA 2 7B 100 71.50 79.89
1000 72.19 72.95
2000 72.18 75.36
Appendix D Model Behavior Analysis
In the reasoning generation tasks, the performance of the model varied across explicit reasoning, comparison reasoning, and their combination. Below, we provide an analysis of unexpected behavior based on 2,000 queries.
Repetition.
The repetition metric measures the occurrence of duplicate passage identifiers. Across the tasks, we observed:
  • Explicit reasoning: 63 duplicate passage identifiers (3%).
  • Comparison reasoning: 36 duplicate passage identifiers (1.5%).
  • Combined reasoning: 47 duplicate passage identifiers (2%).
Missing Documents.
Missing documents were most frequent in explicit reasoning:
  • Explicit reasoning: 642 missing documents (32% of queries).
  • Comparison reasoning: 227 missing documents (11% of queries).
  • Combined reasoning: 328 missing documents (16% of queries).
Appendix E API Cost Analysis
Utilizing large language models (LLMs) like GPT-4 as teacher models involves significant computational and monetary costs. In this section, we analyze these costs based on our experiments.
We calculated the average number of tokens, the number of API requests, and the estimated cost per query for different operations involving the teacher model. The costs are based on the pricing provided by the API service at the time of our experiments 9.
API Instruction Tokens Cost (\$USD)
gpt-4 Explicit 3650 0.134
gpt-4 Comparison 4050 0.158
gpt-4 Comparison & Explicit 4650 0.194
Table 9: Average token usage and cost per query using GPT-4 for different reasoning tasks.
Appendix F Prompts for Reasoning
Table 10: Basic prompt, explicit reasoning prompt, and comparison reasoning prompt.
Basic Prompt:
I will provide you with num passages, each indicated by a number identifier [].
For each passage, briefly generate your reasoning process as follows:
1. Judge whether the passage is not applicable (not very common), where the query does not meet the
premise of the passage.
2. Check if the query contains direct evidence. If so, judge whether the query meets or does not meet
the passage.
3. If there is no direct evidence, try to infer from existing evidence and answer one question:
If the passage is ranked in this order, is it possible that a good passage will miss such information?
If impossible, then you can assume that the passage should not be ranked in that order.
Otherwise, it should be ranked in that order.
Explicit Reasoning Prompt:
Then, read every passage one-by-one and find the sentence where the method is the direct answer
for the query and output the sentence. Then, rank the passages based on their relevance to the query.
Give highest priority to passages that provide a clear and direct definition or explanation of the
keywords related to the query.
Consider both the detailed information and any relevant background context provided in each passage.
Provide clear and concise reasons for the ranking, highlighting the specific parts of the passages that
influenced your decision.
Make sure to ignore any irrelevant information and focus on content directly related to the query.
In addition to direct reasons for ranking each passage, also consider listwise reasons.
Comparison Reasoning Prompt:
A listwise reason involves comparing passages with each other to determine their relative importance.
Specifically, compare passages to identify which ones provide more comprehensive, relevant, and accurate
information in relation to the query.
When providing listwise reasons, mention specific comparative insights, such as why one passage might be
more relevant than another based on the overall context and detail provided.
Return Type:
Search Query: query. Rank the num passages above based on their relevance to the search query and
the extracted keywords. The passages should be listed in descending order using identifiers.
The most relevant passages should be listed first. The output should be in JSON format.
For each passage, provide the extracted keywords and detailed reasons for its ranking.
Ensure to separate direct reasons and listwise reasons clearly.
Mention specific parts of the passage that influenced your decision.
Ensure the reasons are clear, concise, and directly related to the query,
balancing direct definitions or explanations and relevant background information.
Only output the JSON structured format.

### v2

@@
-ReasoningRank: Teaching Student Models to Rank through Reasoning-Based Knowledge Distillation
+Reason-to-Rank: Learning to Rank through Reasoning-Based Knowledge Distillation
 Abstract
 Reranking documents based on their relevance to a given query is a critical task in information retrieval.
-Traditional reranking methods often focus on improving the initial rankings but tend to lack transparency, failing to explain why one document is ranked higher than another.
-In this paper, we introduce ReasoningRank, a novel reranking approach that enhances clarity by generating two types of reasoning: explicit reasoning, which explains how a document addresses the query, and comparison reasoning, which justifies the relevance of one document over another.
-We leverage large language models (LLMs) as teacher models to generate these explanations and distill this knowledge into smaller, more resource-efficient student models.
-While the student models may not outperform LLMs in speed, they significantly reduce the computational burden by requiring fewer resources, making them more suitable for large-scale or resource-constrained settings.
-These student models are trained to both generate meaningful reasoning and rerank documents, achieving competitive performance across multiple datasets, including MSMARCO and BRIGHT.
-Experiments demonstrate that ReasoningRank improves reranking accuracy and provides valuable insights into the decision-making process, offering a structured and interpretable solution for reranking tasks.
+Traditional reranking methods often lack transparency and rely on proprietary models, hindering reproducibility and interpretability.
+We propose Reason-to-Rank (R2R), a novel open-source reranking approach that enhances transparency by generating two types of reasoning: direct relevance reasoning, which explains how a document addresses the query, and comparison reasoning, which justifies the relevance of one document over another.
+We leverage large language models (LLMs) as teacher models to generate these explanations and distill this knowledge into smaller, openly available student models.
+Our student models are trained to generate meaningful reasoning and rerank documents, achieving competitive performance across multiple datasets, including MSMARCO and BRIGHT.
+Experiments demonstrate that R2R not only improves reranking accuracy but also provides valuable insights into the decision-making process.
+By offering a structured and interpretable solution with openly accessible resources, R2R aims to bridge the gap between effectiveness and transparency in information retrieval, fostering reproducibility and further research in the field.
 [^1] Our code and data are available here https://anonymous.4open.science/r/Distillation-Is-All-You-Need-C98D/README.md
-ReasoningRank: Teaching Student Models to Rank through Reasoning-Based Knowledge Distillation
+Reason-to-Rank: Learning to Rank through Reasoning-Based Knowledge Distillation
 Yuelyu Ji1, Zhuochun Li1, Rui Meng2, Daqing He1 1School of Computing and Information, University of Pittsburgh, 2Salesforce Research {yuj49,zhl163, dah44}@pitt.edu, ruimeng@salesforce.com
@@
 Traditional reranking approaches often improve initial rankings provided by retrieval systems but pay limited attention to the reasoning behind ranking decisions.
-Recent studies have demonstrated that incorporating reasoning into the reranking process can enhance performance and interpretability Pradeep et al.
+Recent studies have revealed that incorporating reasoning into the reranking process can enhance performance and interpretability Pradeep et al.
 (2023b); Liu et al.
@@
 By offering explanations that clarify why certain documents are prioritized over others, these methods help users understand and trust the search results.
-However, most existing reranking models do not distinguish between two key types of reasoning: explicit reasoning, which explains how a document matches a query, and comparative reasoning, which justifies why one document ranks higher than another.
+However, most existing reranking models do not distinguish between two key types of reasoning: direct relevance reasoning, which explains how a document matches a query, and comparative reasoning, which justifies why one document ranks higher than another.
 This limitation can reduce ranking effectiveness, particularly in contexts where a subtlety understanding of document relevance is crucial.
-Figure 1: Illustration of the two types of reasoning: explicit reasoning provides direct answers to the query, while comparison reasoning evaluates the relative relevance between documents.
+Figure 1: Illustration of the two types of reasoning: direct relevance reasoning provides explicit answers to the query, while comparison reasoning evaluates the relative relevance between documents.
 The LLM generates these explanations to enhance the interpretability of the ranking process.
-In this paper, we propose ReasoningRank, a reranking framework that leverages both explicit and comparative reasoning to provide more detailed and transparent document rankings.
-Our approach is based on knowledge distillation, where a large, powerful teacher model (an LLM) generates high-quality explanations and rankings.
-These explanations are then distilled into smaller, more computationally efficient student model that perform the reranking tasks.
+In this paper, we propose Reason-to-Rank (R2R), a reranking framework that leverages both explicit and comparative reasoning to provide more detailed and transparent document rankings.
+Our approach based on knowledge distillation, where a large, powerful teacher model (an LLM) generates high-quality explanations and rankings.
+Distilling these explanations into smaller, more computationally efficient student model that perform the reranking tasks.
 Our contributions are threefold:
-  • We introduce a reranking approach that separates explicit and comparative reasoning, enhancing clarity of document rankings.
-  • We demonstrate that student models can learn from LLM-generated data, achieving high reranking performance while remaining computationally efficient.
-  • We provide an open-source implementation to facilitate further research and applications in information retrieval.
-The remainder of this paper is organized as follows.
-Section 2 reviews related work.
-Section 3 presents the proposed ReasoningRank framework.
-Section 4 not only presents the quantitative performance metrics but also provides a thorough discussion on the implications of the results, analyzing how the reasoning strategies impact the ranking performance and offering insights into the model’s interpretability.
-Finally, Section 6 concludes the paper and discusses future work.
+  • Introduce a reranking approach that separates direct relevance and comparative reasoning, enhancing the clarity of document rankings.
+  • Demonstrate that student models can learn from LLM-generated data, achieving high reranking performance while remaining computationally efficient.
+  • Provide an open-source implementation to facilitate further research and applications in information retrieval.
+The remainder of this paper is organize as follows.
+§2 reviews related work.
+§3 presents the Reason-to-Rank (R2R) framework.
+Sections 4 not only present the quantitative performance metrics but also provide a thorough discussion on the implications of the results, analyzing how the reasoning strategies impact the ranking performance and offering insights into the model’s interpretability.
+Finally, §6 concludes the paper and discusses future work.
 2 Related Work
@@
 However, these models rely heavily on large-scale fine-tuning and task-specific data, limiting their scalability and adaptability to new domains or tasks.
-Figure 2: Overview of explicit and comparison reasoning prompts for document ranking.
-Explicit reasoning explains how a document matches a query, while comparison reasoning evaluates the relative relevance between documents.
-Figure 3: Overview of the ReasoningRank framework.
-The teacher model generates explicit and comparative reasoning, which is used to train student models capable of reproducing the reranked order and generating explanations.
-Early ranking systems primarily focused on pointwise methods, where each document’s relevance was judged in isolation Liu et al.
+Figure 2: Overview of direct relevance and comparison reasoning prompts for document ranking.
+Pairwise reasoning explains how a document matches a query, while comparison reasoning evaluates the relative relevance between documents.
+Figure 3: Overview of the Reason-to-Rank framework.
+The teacher model generates direct relevance and comparative reasoning, which is used to train student models capable of reproducing the reranked order and generating explanations.
+Early ranking systems primarily focused on direct relevance methods, where each document’s relevance was judged in isolation Liu et al.
 (2009).
@@
 (2009).
-Pointwise approaches focus on predicting the relevance score of individual documents in isolation, while pairwise methods aim to indicate the relative ranking between document pairs.
+Direct relevance approaches focus on predicting the relevance score of individual documents in isolation, while pairwise methods aim to indicate the relative ranking between document pairs.
 On the other hand, listwise approaches focus on directly optimizing the entire ranking list, offering better overall accuracy by considering the global structure of the ranking.
@@
 (2023) have focused on generating explanations that describe how a document matches a query.
-However, these approaches often only address explicit reasoning (i.e., document-query matching) and do not account for comparative rationale (i.e., why one document ranks higher).
+However, these approaches often only address direct relevance reasoning (i.e., document-query matching) and do not account for comparative rationale (i.e., why one document ranks higher).
 Our work builds on these efforts by incorporating both types of reasoning, enhancing the overall clarity and coherence of the explanations.
@@
 (2015) to further reduce the computational demands of large reranking models.
-Knowledge distillation has been particularly effective in reranking tasks, as demonstrated by models like RankVicuna Pradeep et al.
+Knowledge distillation has been particularly effective in reranking tasks, as shown by models like RankVicuna Pradeep et al.
 (2023a), which distilled the reranking capabilities of RankGPT into a smaller, more efficient model without significant loss of performance.
@@
 (2024b, a).
-By integrating these advancements, our ReasoningRank framework improves upon existing methods by offering a dual reasoning strategy—explicit and comparative reasoning—within a computationally efficient model.
+By integrating these advances, our R2R framework improves existing methods by offering a dual reasoning strategy - direct relevance and comparative reasoning - within a computationally efficient model.
 Our approach not only enhances ranking accuracy but also provides clear and easy-to-understand explanations for the ranking decisions, addressing fundamental limitations in performance seen in prior work.
 3 Method
-The overview of the architecture of the ReasoningRank model is shown in Figure 3.
-We propose ReasoningRank, a framework that integrates explicit and comparative reasoning within a unified reranking approach.
+The overview of the architecture of the R2R model is shown in Figure 3. R2R, a framework that integrates direct relevance and comparative reasoning within a unified reranking approach.
 3.1 Teacher Model with Dual Reasoning
 Given a query $q$ and a set of candidate documents $\{d_{1},d_{2},\dots,d_{n}\}$, the teacher model processes these inputs to generate two types of reasoning and rerank order.
-The teacher model can generate high-quality reasoning that combines both explicit and comparative aspects, and the examples can be seen in Figure 2: Explicit Reasoning: The model explains how each document directly addresses the query, focusing on the relevance and specificity of the content.
+The teacher model can generate high-quality reasoning that combines both direct relevance and comparative aspects, and the examples can be seen in Figure 2: Direct relevance Reasoning: The model explains how each document directly addresses the query, focusing on the relevance and specificity of the content.
 Comparative Reasoning: The model assesses the relative relevance between documents, explaining why one document should be ranked higher based on content.
@@
 Deploying large language models (LLMs) for reranking is often impractical due to their high computational cost.
-To address this, we designed the student model to replicate the teacher’s performance while being more efficient for real-world deployment.
-The innovation lies in using knowledge distillation: the student learns from the teacher’s detailed reasoning and ranking outputs, retaining accuracy and interpretability with significantly lower computational demands.
+To address this, we design the student model to replicate the teacher’s performance while being more efficient for real-world deployment.
+The innovation lies in using knowledge distillation: the student learns from the teacher’s reasoning and ranking outputs, retaining accuracy and interpretability with significantly lower computational demands.
 This approach ensures the student can perform complex reranking tasks efficiently, making it ideal for large-scale use.
-Architecture and Training:
+3.2.1 Model Overview
 Given a query $q$ and a set of candidate documents $\{d_{1},d_{2},\dots,d_{n}\}$, the model processes these inputs to generate both reasoning outputs and ranking scores.
@@
 Ranking Scores: The model simultaneously generates ranking scores $\mathbf{s}_{i}$ for each document, indicating its relevance to the query.
-3.2.1 Training Objective and Loss Functions
+3.2.2 Training Objective
 To optimize the performance of the student model, we define three distinct loss functions, each corresponding to different aspects of the model.
-First, using pairwise loss to optimize the relative ranking between pairs of documents, ensuring that the more relevant document is ranked higher.
+First, the pairwise loss optimizes the relative ranking between pairs of documents, ensuring that the more relevant document is ranked higher.
 The pairwise loss is defined as:
@@
 Next, the listwise loss optimizes the overall ranking of the document list.
-It uses Kullback-Leibler (KL) divergence to measure the difference between the predicted ranking distribution and the target ranking distribution:
+It uses Kullback-Leibler (KL) divergence Hershey and Olsen (2007) to measure the difference between the predicted ranking distribution and the target ranking distribution:
 $\mathcal{L}_{\text{listwise}}=D_{\text{KL}}(Q(\mathbf{z})\parallel P(\mathbf{s}))$
 where $P(\mathbf{s})=\text{softmax}(\mathbf{s}_{i})$ is the predicted ranking distribution based on the model’s output, and $Q(\mathbf{z})=\text{softmax}(\mathbf{z}_{i})$ is the target ranking distribution based on true labels.
-Finally, the generation loss optimizes the quality of the textual reasoning generated by the model. Based on the cross-entropy loss between the generated reasoning and the ground-truth reasoning:
+Finally, the generation loss optimizes the quality of the textual reasoning generated by the model.
+It is based on the cross-entropy loss between the generated reasoning and the ground-truth reasoning:
 $\mathcal{L}_{\text{generation}}=-\sum_{i=1}^{n}\mathbf{y}_{i}^{\text{gen}}\log(\text{softmax}(\mathbf{r}_{i}))$
-where $\mathbf{r}_{i}$ represents the generated reasoning logits for document $d_{i}$, and $\mathbf{y}_{i}^{\text{gen}}$ is the ground truth reasoning label.
-3.3 Final Training Objective
+where $\mathbf{r}_{i}$ represents the generated reasoning logits for document $d_{i}$, and $\mathbf{y}_{i}^{\text{gen}}$ is the ground-truth reasoning label.
 To achieve balanced training, we introduce learnable weights $\alpha$, $\beta$, and $\gamma$ to adjust the contributions of pairwise, listwise, and generation losses, respectively.
@@
 With this design, the student model efficiently performs ranking tasks and provides transparent reasoning for its decisions, enhancing both transparency and performance.
-4 Experiments and Results
-In this section, we evaluate the performance of our proposed ReasoningRank framework across various information retrieval tasks.
-We introduce the datasets and evaluation metrics used in our experiments (§4.1).
-We then detail the experimental setup and baseline methods for comparison (§4.3).
-Finally, we present a comprehensive analysis of the experimental results, including comparisons with existing methods in (§4.4).
-We performed paired t-tests across all queries in each dataset to assess the statistical significance of the improvements.
-The annotations in Tables 1 and 2 indicate the significance levels.
+Models DL19 DL20 BEIR-6 Avg. Covid Touche News NFCorpus Robust04 DBPedia
+Baseline Models
+DeBERTa Sun et al. (2023) 68.5 64.2 47.0 73.4 32.1 50.2 33.7 49.2 45.4
+MonoT5 Nogueira et al. (2020) 74.5 70.4 48.3 80.0 34.1 46.0 35.2
+RankVicuna Pradeep et al. (2023a) 71.1 68.7 48.3 67.1 48.7 38.5 55.7 35.3
+RankZephyr Pradeep et al. (2023b) 72.2 70.5 51.3 85.1 36.5 53.3 38.9 60.7 35.5
+APEER Jin et al. (2024) 74.6 72.3 51.1 83.9 35.3 52.1 33.4 56.0 46.1
+Our Models (student)
+w/o reasoning (Pairwise Loss) 73.2 70.8 50.1 79.9 35.1 51.9 34.6 52.8 46.3
+w/o reasoning (Listwise Loss) 73.8 71.0 50.5 80.2 35.4 52.3 35.0 53.0 46.6
+w/ direct relevance reasoning 74.5 70.1 49.8 79.4 34.3 52.8 35.5 50.8 47.0
+w/ comparison reasoning 74.1 72.3 50.9 80.1 34.0 53.1 36.6 52.1 47.2
+w/ direct relevance & comparison 75.4 72.4 52.4 84.6 36.2 53.8 36.4 53.5 47.9
+Table 1: NDCG@5 performance (in percentage) for student models and baseline comparisons across multiple datasets.
+’w/o Reasoning’ models do not use reasoning prompts and do not generate explanations; they are optimized using the specified ranking loss functions.
+’w/ direct relevance Reasoning’ and ’w/ comparison Reasoning’ models use reasoning prompts and include reasoning loss for generating explanations.
+’’ indicates statistically significant improvement over the baseline DeBERTa model (p $<$ 0.05).
+Bold indicates the best performance among all models for each dataset.
+Entries marked with ’–’ indicate that results are not available for that dataset.
+4 Experiments
+In this section, we evaluate the performance of our Reason-to-Rank (R2R) framework across various information retrieval tasks.
+We introduce the datasets and evaluation metrics used in our experiments §4.1.
+We then detail the experimental setup and baseline methods for comparison §4.2.
+Finally, we present a comprehensive analysis of the experimental results, including comparisons with existing methods in §4.2.4 and §4.2.1.
 4.1 Datasets and Evaluation Metrics
@@
 Normalized Discounted Cumulative Gain (NDCG@k): Measures the ranking quality of the retrieved documents, with higher emphasis on top-ranked documents. We report NDCG at rank positions 5 and 10.
-BLEU, ROUGE-L, and Coherence Score: Used for evaluating the quality of generated reasons in reasoning tasks, assessing both lexical overlap and logical consistency.
-4.2 Teacher Models
-We utilize large language models (LLMs) as teacher models to generate reasoning annotations: GPT-4: A state-of-the-art LLM known for its strong reasoning and language understanding capabilities.
-ClaudeThe : An advanced LLM developed by Anthropic, offering competitive performance in reasoning tasks, we used the version claude-3-5-sonnet-20240620.
-GeminiTeam et al.
-(2023): A commercial LLM used for comparison purposes.We used the version gemini-1.5-flash.
-4.2.1 Overall Performance on Benchmark Datasets
-Table 1 summarizes the NDCG@5 performance of different models GPT-4 consistently achieves the best results, with 77.69% on DL19 and 85.27% on TREC-COVID when using both comparison and explicit reasoning prompts, showcasing its robust handling of complex queries.
+BLEU, and ROUGE-L: Used for evaluating the quality of generated reasons in reasoning tasks, assessing both lexical overlap and logical consistency.
+4.2 Models
+4.2.1 Teacher models
+We utilize large language models (LLMs) as teacher models to generate reasoning annotations: GPT-4 Achiam et al.
+(2023), Claude The (claude-3-5-sonnet-20240620) and Gemini Team et al.
+(2023) (gemini-1.5-flash).
+These advanced LLMs known for their strong reasoning and language-understanding capabilities.
+Teacher Model Performance
+Table 5 summarizes the NDCG@5 performance of different models GPT-4 consistently achieves the best results, with 77.6% on DL19 and 85.2% on TREC-COVID when using both comparison and direct relevance reasoning prompts, showcasing its robust handling of complex queries.
 Adding different reasoning prompts to the LLMs enhances their performance across various datasets.
-For instance, GPT-4 with comparison and explicit reasoning outperforms its baseline (without reasoning) on DL19 by 2.10%.
-Similar trends are observed with Claude and Gemini, indicating that incorporating reasoning prompts in different LLMs is beneficial.
-To confirm the statistical significance of the improvements, we conducted paired t-tests comparing models with reasoning prompts to their respective baselines.
-The results, indicated by and in Table 1, show that the enhancements are significant, with p-values less than 0.05.
+For instance, GPT-4 with comparison and direct relevance reasoning outperforms its baseline (without reasoning) on DL19 by 2.1%.
+Similar trends are observesd with Claude and Gemini, indicating that incorporating reasoning prompts in different LLMs is beneficial.
+To confirm the statistical significance of the improvements, we conduct paired t-tests comparing models with reasoning prompts to their respective baselines.
 Given GPT-4’s overall solid performance and the boost from reasoning prompts, we select it as the teacher model to distill its enhanced reasoning abilities into our student models.
-Models DL19 DL20 Covid Touche News NFCorpus Robust04 DBPedia
-BM25 50.58 47.96 59.47 44.22 39.52 30.75 40.73 31.83
-Rank GPT-4 Models
-Rank GPT-4 (without reasoning) 75.59 70.56 85.51 38.47 57.55 36.62 57.55 47.12
-+ Explicit reasoning 75.32 69.93 83.27 38.94 55.87 36.08 52.89 46.98
-+ Comparison reasoning 76.25 71.07 84.62 38.23 56.49 36.52 56.53 48.15
-+ Comparison & Explicit 77.69 73.24 85.27 38.29 58.43 36.33 58.63 49.52
-Claude Models
-Claude (without reasoning) 71.96 68.81 82.77 36.12 54.35 36.36 50.78 42.51
-+ Explicit reasoning 70.59 68.23 82.22 35.92 55.43 35.59 52.38 42.51
-+ Comparison reasoning 71.34 69.75 83.31 36.43 55.11 36.71 51.53 43.23
-+ Comparison & Explicit 72.13 69.95 84.03 37.33 55.52 36.36 52.67 44.91
-Gemini Models
-Gemini (without reasoning) 71.10 68.45 82.95 35.62 53.65 35.49 50.12 42.92
-+ Explicit reasoning 69.08 67.02 81.55 34.83 52.87 36.92 51.31 43.92
-+ Comparison reasoning 70.17 68.24 82.41 35.91 53.35 35.42 51.12 42.85
-+ Comparison & Explicit 71.35 68.57 83.47 36.97 53.80 36.12 51.78 43.67
-Table 1: NDCG@5 performance (in percentage) for different models with various reasoning prompts across multiple datasets.
-Bold numbers indicate the best performance within each group of models.
-indicates statistically significant improvement over the base model without reasoning prompts (p $<$ 0.05).
-indicates statistically significant improvement over both the base model and the single reasoning prompt variants (p $<$ 0.05).
-Models DL19 DL20 Covid Touche News NFCorpus Robust04 DBPedia
-Student Models
-Basic w/o reasoning 73.23 70.80 79.94 35.10 51.90 34.60 52.82 46.31
-With explicit reasoning 74.54 70.11 79.37 34.25 52.81 35.45 50.75 46.95
-With comparison reasoning 74.12 72.25 80.10 34.00 53.10 36.63 52.11 47.21
-Comparison & Explicit 75.39 72.36 84.59 36.22 53.81 36.37 53.54 47.85
-Baseline Models
-DeBERTa Sun et al. (2023) 68.53 64.18 73.42 32.13 50.24 33.67 49.23 45.39
-MonoT5 Nogueira et al. (2020) 74.54 70.44 79.95 34.07 46.03 35.21
-RankVicuna Pradeep et al. (2023a) 71.12 68.66 67.14 48.73 38.53 55.67 35.29
-RankZephyr Pradeep et al. (2023b) 72.23 70.53 85.12 36.46 53.33 38.92 60.72 35.52
-APEER Jin et al. (2024) 74.57 72.25 83.86 35.25 52.11 33.41 56.03 46.13
-Table 2: NDCG@5 performance (in percentage) for student models and baseline comparisons across multiple datasets.
-indicates statistically significant improvement over the baseline DeBERTa model (p $<$ 0.05).
-Bold indicates the best performance among all models for each dataset.
-Entries marked with "–" indicate that results are not available for that dataset.
-4.3 Student Models
-Experimental Setup
+4.2.2 Student models
 We implement our models using LoRA (Low-Rank Adaptation) on 32GB V100 GPUs to fine-tune the student models on large datasets efficiently.
 Based on the LLaMA 3.1 8B architecture, the student models fine-tuned with different reasoning prompts derived from the teacher models.
-4.3.1 Performance Comparison with Baseline
-Our student models, trained with explicit and comparison reasoning, achieve competitive performance compared to the previous state-of-the-art student models.
-As shown in Table 2,the improvements over the baseline models are statistically significant, with p-values less than 0.05 (indicated by ).
+Impact of Reasoning Strategies
+we also evaluated different reasoning strategies in the R2R student models, specifically analyzing the "w/o Reasoning" (without reasoning), "w/ direct relevance reasoning" (with direct relevance reasoning), "w/ comparison reasoning" (with comparison reasoning), and "w/ direct relevance & comparison" (with both direct relevance and comparison reasoning) variants.
+As shown in Table  1, models without reasoning (using pairwise or listwise loss) perform at a baseline level, with NDCG@5 scores of 73.2% and 73.8 on DL19 and 70.8% and 71.0% on DL20.
+Adding reasoning significantly boosts performance.
+For example, direct relevance reasoning improves NDCG@5 to 74.5% on DL19 and 70.1% on DL20 by providing direct relevance relevance explanations.
+Comparison reasoning, which assesses the relative importance between documents, raises NDCG@5 to 74.1% on DL19 and 72.3% on DL20.
+Combining both types of reasoning yields the best results, with NDCG@5 scores of 75.4% on DL19 and 72.4% on DL20.
+Interestingly, while reasoning improves performance in most cases, there are instances where models without reasoning (e.g., listwise loss) outperform those with reasoning.
+4.2.3 Comparison with Baselines
+Our student models, trained with direct relevance and comparison reasoning, achieve competitive performance compared to the previous state-of-the-art student models.
+As shown in Table  1,the improvements over the baseline models are statistically significant, with p-values less than 0.05.
 BM25: A traditional term-based retrieval model using TF-IDF, known for its simplicity and efficiency in keyword-based retrieval.
@@
 (2023a, b); Jin et al.
-(2024): RankVicuna is a distilled version of RankGPT, designed for zero-shot listwise reranking with fewer parameters, balancing efficiency and performance.RankZephyr a state-of-the-art zero-shot listwise reranking model that adapts RankGPT prompts for robust performance with minimal task-specific training.
+(2024): RankVicuna is a distilled version of RankGPT, designed for zero-shot listwise reranking with fewer parameters, balancing efficiency and performance.
+RankZephyr a state-of-the-art zero-shot listwise reranking model that adapts RankGPT prompts for robust performance with minimal task-specific training.
 APEER share the same structure with RankZephyr but have an automated prompt generation for large language models.
-4.4 ReasoningRank Student Model vs. Baseliness
-Our ReasoningRank student model demonstrates strong performance across various datasets, often outperforming the baseline models.
-For instance, on the DL19 dataset, our student model achieves an NDCG@5 of 75.39%, surpassing DeBERTa (68.53%), RankVicuna (71.12%), and RankZephyr (72.23%).
-Similarly, on the DL20 dataset, our model attains 72.36%, outperforming DeBERTa (64.18%) and RankVicuna (68.66%).
+4.2.4 R2R Student Model vs. Baselines
+Our R2R student model indicates strong performance across various datasets, often outperforming the baseline models.
+For instance, on the DL19 dataset, our student model achieves an NDCG@5 of 75.3%, surpassing DeBERTa (68.5%), RankVicuna (71.1%), and RankZephyr (72.2%).
+Similarly, on the DL20 dataset, our model attains 72.3%, outperforming DeBERTa (64.1%) and RankVicuna (68.6%).
 On some datasets, our model achieves the best performance among all models.
-For example, on the News dataset, our student model reaches an NDCG@5 of 53.81%, slightly outperforming RankZephyr (53.33%) and significantly better than DeBERTa (50.24%).
+For example, on the News dataset, our student model reaches an NDCG@5 of 53.8%, slightly outperforming RankZephyr (53.3%) and significantly better than DeBERTa (50.2%).
 However, there are cases where other models perform better.
-On the Touche dataset, RankVicuna achieves an NDCG@5 of 48.73%, which is higher than our student model’s 36.22%.
-This may be due to RankVicuna’s design for zero-shot listwise reranking, which could be particularly effective for argument retrieval tasks represented in Touche.
-Similarly, on the Robust04 dataset, RankZephyr attains a higher NDCG@5 of 60.72% compared to our model’s 53.50%.
-These variations suggest that while our model performs well, certain models may have advantages in specific domains or datasets due to their architectures or training strategies.
-For example, RankVicuna’s strong performance on Touche might be attributed to its effectiveness in handling argumentation and persuasive texts, which are prevalent in that dataset.
-RankZephyr’s higher score on Robust04 could be due to its ability to leverage zero-shot capabilities effectively in that particular dataset.
+On the Touche dataset, RankVicuna achieves an NDCG@5 of 48.7%, which is higher than our student model’s 36.2%.
+This may be due to RankVicuna’s design for zero-shot listwise reranking, which could be particularly effective for argument retrieval tasks represent in Touche.
+Similarly, on the Robust04 dataset, RankZephyr attains a higher NDCG@5 of 60.7% compared to our model’s 53.5%.
 4.2.5 Performance on Reasoning-Intensive Tasks (BRIGHT Dataset)
-The BRIGHT dataset is specifically designed to evaluate models on reasoning-intensive retrieval tasks, simulating real-world scenarios where complex reasoning is required to determine document relevance.
+The BRIGHT dataset Su et al.
+(2024) is specifically designed to evaluate models on reasoning-intensive retrieval tasks, simulating real-world scenarios where complex reasoning is required to determine document relevance.
+We report the Inst-LSu et al.
+(2024, 2022) model as a baseline because it is fine-tuned with task-specific instructions, making it particularly well-suited for tasks requiring deeper reasoning.
 Table 2 shows the NDCG@5 results on the BRIGHT dataset.
-In the Earth Science domain, the student model achieves an NDCG@5 of 27.0%, closely approaching the teacher model’s 31.0%.
-Similarly, in Psychology, the student model attains 26.0%, compared to the teacher’s 28.9% and BM25’s 11.6%.
-Domain BM25 Teacher Student GPT-4 COT
-Bio. 17.2 24.3 21.0 28.0
-Earth. 24.2 31.0 27.0 37.1
-Econ. 13.2 19.0 16.5 22.1
-Psy. 11.6 28.9 26.0 34.2
-Rob. 12.0 16.0 14.5 19.0
-Stack 16.0 21.0 18.5 24.2
-Sus. 12.4 18.0 16.2 21.9
-Leet. 22.1 27.6 25.0 31.4
-Pony 8.5 5.9 5.2 6.7
-AoPS 6.0 7.6 6.8 9.2
-TheoQ. 9.4 18.5 15.3 23.0
-TheoT. 3.5 17.0 14.2 21.7
-Average 13.1 19.6 17.5 23.2
+In the Earth Science domain, our R2R model achieves an NDCG@5 of 31.0%, closely approaching the teacher model’s 37.1%.
+Similarly, in Psychology, the student model attains 28.9%, compared to the teacher’s 34.2% and BM25’s 11.6%.
+However, we observe that in certain domains, such as Pony and AoPS, the performance of our student and teacher models is relatively low.
+Domain BM25 Inst-L Su et al. (2024) R2R GPT-4 Teacher
+Bio. 17.2 21.0 24.3 28.0
+Earth. 24.2 27.0 31.0 37.1
+Econ. 13.2 16.5 19.0 22.1
+Psy. 11.6 26.0 28.9 34.2
+Rob. 12.0 14.5 16.0 19.0
+Stack 16.0 18.5 21.0 24.2
+Sus. 12.4 16.2 18.0 21.9
+Leet. 22.1 25.0 27.6 31.4
+Pony 8.5 5.2 5.9 6.7
+AoPS 6.0 6.8 7.6 9.2
+TheoQ. 9.4 15.3 18.5 23.0
+TheoT. 3.5 14.2 17.0 21.7
+Average 13.1 17.5 19.6 23.2
 Table 2: NDCG@5 results on the BRIGHT dataset.
-4.4.2 Quality of Generated Reasons
+5 Ablation Study
+5.1 Quality of Generated Reasons
 We assess the quality of the generated reasons using BLEU and ROUGE-L as presented in Table 3 compared with the teacher model’s reason.
@@
 Similar improvements are observed on DL20, where the student model attains BLEU and ROUGE-L scores of 24.0 and 38.2, respectively, surpassing the baseline’s 19.5 and 34.0.
-Dataset Model BLEU ROUGE-L
-DL19 Baseline 18.0 32.5
-Ours 21.4 36.8
-DL20 Baseline 19.5 34.0
-Ours 24.0 38.2
-Covid Baseline 16.0 30.0
-Ours 18.9 34.5
-Touche Baseline 17.5 33.0
-Ours 20.5 37.1
-News Baseline 20.0 35.0
-Ours 22.7 39.3
-Table 4: Evaluation of reason generation using BLEU, and ROUGE-L scores for baseline and student models.
-4.4.3 Ablation Study: Impact of Reasoning Strategies
-Results
-In Table 5, by removing comparison reasoning results in a performance drop on DL19 from 75.39 to 74.12 (NDCG@5).
-Similarly, removing explicit reasoning causes a performance decrease on DL20 from 72.36 to 72.25 (NDCG@5).
-Analysis
+Dataset Student Model BLEU ROUGE-L
+DL19 w/o Rank Loss 18.0 32.5
+w/ Rank Loss 21.4 36.8
+DL20 w/o Rank Loss 19.5 34.0
+w/ Rank Loss 24.0 38.2
+Covid w/o Rank Loss 16.0 30.0
+w/ Rank Loss 18.9 34.5
+Touche w/o Rank Loss 17.5 33.0
+w/ Rank Loss 20.5 37.1
+News w/o Rank Loss 20.0 35.0
+w/ Rank Loss 22.7 39.3
+Table 3: Evaluation of reason generation using BLEU and ROUGE-L scores for student models trained with and without the ranking loss function.
+The ’Student Model w/o Rank Loss’ is trained without the rank loss function and has the same input as the ’Student Model w/ Rank Loss’.
+5.2 Impact of Reasoning Strategies
+In Table 4, by removing comparison reasoning results in a performance drop on DL19 from 75.3 to 74.1 (NDCG@5).
+Similarly, removing direct relevance reasoning causes a performance decrease on DL20 from 72.3 to 72.2 (NDCG@5).
 Both reasoning strategies contribute to performance improvements.
-The ablation study shows that omitting either reasoning component results in a noticeable decline in performance, highlighting their complementary roles in improving ranking effectiveness.
 Model DL19 NDCG@5/@10 DL20 NDCG@5/@10
-Comparison & explicit 75.39 / 73.85 72.36 / 71.22
-w/ explicit reasoning 74.54 / 73.18 70.11 / 70.97
-w/ comparison reasoning 74.12 / 72.94 72.25 / 70.43
-Basic w/o reasoning 73.23 / 72.12 70.82 / 69.83
-Table 5: Ablation study: Impact of explicit and comparison reasoning on NDCG@5/@10 for DL19 and DL20.
-4.4.4 Effects of Training Data Size Across Student Models
+Direct relevance & comparison 75.3 / 73.8 72.3 / 71.22
+w/ direct relevance reasoning 74.5 / 73.1 70.1 / 70.9
+w/ comparison reasoning 74.1 / 72.9 72.2 / 70.4
+w/o reasoning 73.2 / 72.1 70.8 / 69.8
+Table 4: Ablation study: Impact of direct relevance and comparison reasoning on NDCG@5/@10 for DL19 and DL20.
+5.3 Effects of Training Data Size
 Figure 4 i shows the performance of small models trained on 100, 1,000, and 2,000 examples.
 Contrary to expectations, increasing data size from 1,000 to 2,000 does not consistently improve performance and sometimes slightly reduces it.
-We observed similar trends across various small models, including Mistral and LLaMA variants, highlighting that adding more data does not always guarantee better performance.
+We find similar trends across various small models, including Mistral and LLaMA variants, highlighting that adding more data does not always guarantee better performance.
 Figure 4: Student model performance for different training data sizes on DL19 and DL20.
+Models DL19 DL20 BEIR-6 Avg. Covid Touche News NFCorpus Robust04 DBPedia
+BM25 50.6 48.0 41.0 59.5 44.2 39.5 30.8 40.7 31.8
+Rank GPT-4 Model
+Rank GPT-4 (w/o reasoning) 75.6 70.6 54.0 85.5 38.5 57.6 36.6 57.6 47.1
++ pairwise reasoning 75.3 69.9 52.9 83.3 38.9 55.9 36.1 52.9 47.0
++ comparison reasoning 76.3 71.1 53.4 84.6 38.2 56.5 36.5 56.5 48.2
++ direct relevance & comparison 77.7 73.2 54.4 85.3 38.3 58.4 36.3 58.6 49.5
+Claude Model
+Claude (w/o reasoning) 72.0 68.8 50.8 82.8 36.1 54.4 36.4 50.8 42.5
++ pairwise reasoning 70.6 68.2 50.3 82.2 35.9 55.4 35.6 52.4 42.5
++ comparison reasoning 71.3 69.8 51.1 83.3 36.4 55.1 36.7 51.5 43.2
++ direct relevance & comparison 72.1 70.0 52.0 84.0 37.3 55.5 36.4 52.7 44.9
+Gemini Model
+Gemini (w/o reasoning) 71.1 68.5 50.1 83.0 35.6 53.7 35.5 50.1 42.9
++ pairwise reasoning 69.1 67.0 51.1 81.6 34.8 52.9 36.9 51.3 43.9
++ comparison reasoning 70.2 68.2 50.2 82.4 35.9 53.4 35.4 51.1 42.9
++ direct relevance & comparison 71.4 68.6 51.0 83.5 37.0 53.8 36.1 51.8 43.7
+Table 5: NDCG@5 performance (in percentage) for different models with various reasoning prompts across multiple datasets.
+Bold numbers indicate the best performance within each group of models.
+indicates statistically significant improvement over the base model without reasoning prompts (p $<$ 0.05).
+indicates statistically significant improvement over both the base model and the single reasoning prompt variants (p $<$ 0.05).
 6 Discussion
-Our experiments demonstrate that the ReasoningRank framework effectively enhances document reranking by incorporating explicit and comparative reasoning.
-The student models achieve competitive results with significantly reduced computational resources compared to large teacher models.
-The statistically significant improvements confirm the effectiveness of reasoning strategies in enhancing model performance.
-An important aspect of ReasoningRank is its ability to not only rank documents but also generate coherent and relevant reasoning for its rankings.
-This dual capability suggests that the model could be applied to other tasks that require both ranking and reasoning generation.
-While our current work focuses on document reranking in information retrieval, the approach of integrating reasoning generation with ranking decisions could be extended to other domains.
-For example, in recommendation systems, the model could rank products or services while providing explanations for its recommendations, thereby improving user trust.
-Similarly, in question answering or summarization tasks, generating reasoning alongside answers or summaries could enhance understandabilityJiang et al.
+Our experiments show that the R2R framework effectively enhances document reranking by incorporating direct relevance and comparative reasoning.
+The student models achieve competitive results while significantly reducing computational resources compared to large teacher models.
+The statistically significant improvements confirm that reasoning strategies effectively enhance model performance.
+An important aspect of R2R is its ability to not only rank documents but also generate coherent and relevant reasoning for its rankings.
+This dual capability suggests that the model could be applied to other tasks requiring both ranking and reasoning generation.
+While our current work centers on document reranking in information retrieval, the integration of reasoning with ranking decisions holds promise for broader applications.
+For instance, in recommendation systems, the model could rank products or services while providing explanations for its choices, enhancing user trust.
+Similarly, in question answering or summarization tasks, generating reasoning alongside answers or summaries could improve understandability Jiang et al.
 (2024).
-However, we acknowledge that our paper does not evaluate the model’s generative abilities on tasks beyond document reranking.
-This is a limitation in assessing the full scope of the model’s applicability.
-Future work should explore how the model performs in different contexts to verify its generalizabilityMcElfresh et al.
+However, future work is needed to explore the model’s performance in these varied contexts and verify its generalizability beyond document reranking McElfresh et al.
 (2022); Rahmani et al.
@@
 7 Conclusion
-In this paper, we introduced ReasoningRank, a novel approach to document reranking that incorporates explicit and comparative reasoning.
+In this paper, we introduced ReasoningRank, a novel approach to document reranking that incorporates direct relevance and comparative reasoning.
 Our method leverages the power of large language models to generate interpretable reasons, which are then distilled into smaller, more efficient student models.
-Our experiments demonstrate the effectiveness of this approach, achieving high performance across multiple datasets while maintaining interpretability.
+Our experiments show the effectiveness of this approach, achieving high performance across multiple datasets while maintaining interpretability.
 Limitations
@@
 Appendix D Model Behavior Analysis
-In the reasoning generation tasks, the performance of the model varied across explicit reasoning, comparison reasoning, and their combination.
+In the reasoning generation tasks, the model’s performance varied across direct relevance reasoning, comparison reasoning, and their combination.
 Below, we provide an analysis of unexpected behavior based on 2,000 queries.
 Repetition.
-The repetition metric measures the occurrence of duplicate passage identifiers. Across the tasks, we observed:
-  • Explicit reasoning: 63 duplicate passage identifiers (3%).
+The repetition metric measures the occurrence of duplicate passage identifiers. Across the tasks, we find:
+  • Pairwise reasoning: 63 duplicate passage identifiers (3%).
   • Comparison reasoning: 36 duplicate passage identifiers (1.5%).
@@
 Missing Documents.
-Missing documents were most frequent in explicit reasoning:
-  • Explicit reasoning: 642 missing documents (32% of queries).
+Missing documents were most frequent in direct relevance reasoning:
+  • Pairwise reasoning: 642 missing documents (32% of queries).
   • Comparison reasoning: 227 missing documents (11% of queries).
@@
 API Instruction Tokens Cost (\$USD)
-gpt-4 Explicit 3650 0.134
+gpt-4 Direct relevance 3650 0.134
 gpt-4 Comparison 4050 0.158
-gpt-4 Comparison & Explicit 4650 0.194
+gpt-4 Direct relevance & Comparison 4650 0.194
 Table 9: Average token usage and cost per query using GPT-4 for different reasoning tasks.
 Appendix F Prompts for Reasoning
-Table 10: Basic prompt, explicit reasoning prompt, and comparison reasoning prompt.
+Table 10: Basic prompt, direct relevance reasoning prompt, and comparison reasoning prompt.
 Basic Prompt:
@@
 Otherwise, it should be ranked in that order.
-Explicit Reasoning Prompt:
-Then, read every passage one-by-one and find the sentence where the method is the direct answer
+Direct relevance Reasoning Prompt:
+Then, read every passage one by one and find the sentence where the method is the direct answer
 for the query and output the sentence. Then, rank the passages based on their relevance to the query.

### v4

@@
-Reason-to-Rank: Learning to Rank through Reasoning-Based Knowledge Distillation
-Abstract
-Reranking documents based on their relevance to a given query is a critical task in information retrieval.
-Traditional reranking methods often lack transparency and rely on proprietary models, hindering reproducibility and interpretability.
-We propose Reason-to-Rank (R2R), a novel open-source reranking approach that enhances transparency by generating two types of reasoning: direct relevance reasoning, which explains how a document addresses the query, and comparison reasoning, which justifies the relevance of one document over another.
-We leverage large language models (LLMs) as teacher models to generate these explanations and distill this knowledge into smaller, openly available student models.
-Our student models are trained to generate meaningful reasoning and rerank documents, achieving competitive performance across multiple datasets, including MSMARCO and BRIGHT.
-Experiments demonstrate that R2R not only improves reranking accuracy but also provides valuable insights into the decision-making process.
-By offering a structured and interpretable solution with openly accessible resources, R2R aims to bridge the gap between effectiveness and transparency in information retrieval, fostering reproducibility and further research in the field.
-[^1] Our code and data are available here https://anonymous.4open.science/r/Distillation-Is-All-You-Need-C98D/README.md
-Reason-to-Rank: Learning to Rank through Reasoning-Based Knowledge Distillation
-Yuelyu Ji1, Zhuochun Li1, Rui Meng2, Daqing He1 1School of Computing and Information, University of Pittsburgh, 2Salesforce Research {yuj49,zhl163, dah44}@pitt.edu, ruimeng@salesforce.com
-1 Introduction
-The ability to effectively rank and rerank documents is critical in information retrieval (IR), where the quality of search results significantly impacts user satisfaction Pradeep et al.
-(2023b); Sun et al.
-(2023).
-Traditional reranking approaches often improve initial rankings provided by retrieval systems but pay limited attention to the reasoning behind ranking decisions.
-Recent studies have revealed that incorporating reasoning into the reranking process can enhance performance and interpretability Pradeep et al.
-(2023b); Liu et al.
-(2024); Yu et al.
-(2024c); Sun et al.
-(2023).
-By offering explanations that clarify why certain documents are prioritized over others, these methods help users understand and trust the search results.
-However, most existing reranking models do not distinguish between two key types of reasoning: direct relevance reasoning, which explains how a document matches a query, and comparative reasoning, which justifies why one document ranks higher than another.
-This limitation can reduce ranking effectiveness, particularly in contexts where a subtlety understanding of document relevance is crucial.
-Figure 1: Illustration of the two types of reasoning: direct relevance reasoning provides explicit answers to the query, while comparison reasoning evaluates the relative relevance between documents.
+Reason-to-Rank: Distilling Direct and Comparative Reasoning from Large Language Models for Document Reranking
+Abstract.
+Reranking documents in information retrieval often relies on black-box models that improve effectiveness but lack explainability.
+We introduce Reason-to-Rank (R2R), a novel framework that separates direct relevance reasoning from comparison reasoning to provide both direct and comparitive explanations.
+We first prompt a large language model to produce comprehensive rationales and a ranking order; then we distill both the ranking decisions and textual explanations into a smaller, open-source student model.
+Our approach not only improves retrieval performance, as demonstrated in MSMARCO, BEIR, and BRIGHT, but also provides interpretable justifications for why one document outranks another.
+We report NDCG@5 (and NDCG@10) for direct comparisons with prior work, and show that the distilled student model achieves competitive results while significantly reducing computational overhead.
+By unifying direct and comparative reasoning in a single pipeline, R2R bridges the gap between transparency and effectiveness in modern reranking systems.
+Large Language Models; Information Retrieval; Re-ranking
+journalyear: 2025copyright: ccconference: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 13–18, 2025; Padua, Italybooktitle: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’25), July 13–18, 2025, Padua, Italydoi: 10.1145/3726302.3730070isbn: 979-8-4007-1592-1/2025/07ccs: Information systems Learning to rank
+1. Introduction
+Figure 1.
+Illustration of the two types of reasoning: direct relevance reasoning provides explicit answers to the query, while comparison reasoning evaluates the relative relevance between documents.
 The LLM generates these explanations to enhance the interpretability of the ranking process.
-In this paper, we propose Reason-to-Rank (R2R), a reranking framework that leverages both explicit and comparative reasoning to provide more detailed and transparent document rankings.
-Our approach based on knowledge distillation, where a large, powerful teacher model (an LLM) generates high-quality explanations and rankings.
-Distilling these explanations into smaller, more computationally efficient student model that perform the reranking tasks.
-Our contributions are threefold:
-  • Introduce a reranking approach that separates direct relevance and comparative reasoning, enhancing the clarity of document rankings.
-  • Demonstrate that student models can learn from LLM-generated data, achieving high reranking performance while remaining computationally efficient.
-  • Provide an open-source implementation to facilitate further research and applications in information retrieval.
-The remainder of this paper is organize as follows.
-§2 reviews related work.
-§3 presents the Reason-to-Rank (R2R) framework.
-Sections 4 not only present the quantitative performance metrics but also provide a thorough discussion on the implications of the results, analyzing how the reasoning strategies impact the ranking performance and offering insights into the model’s interpretability.
-Finally, §6 concludes the paper and discusses future work.
-2 Related Work
-The evolution of reranking techniques has significantly advanced with the rise of neural networks, particularly transformers Vaswani (2017); Yu et al.
-(2024b, a); Zheng et al.
-(2024); Zhang et al.
-(2024); Shen et al.
-(2024); Fan and Tao (2024); Fan et al.
-(2024).
-Early models like monoBERT Nogueira et al.
-(2019) utilized BERT Devlin (2018) in a cross-encoder setup, effectively capturing contextual dependencies between queries and documents.
-Further models such as monoT5 Nogueira et al.
-(2020) and InPars Bonifacio et al.
-(2022) continued to push reranking performance boundaries, especially on datasets like MS MARCO Bajaj et al.
-(2016).
-However, these models rely heavily on large-scale fine-tuning and task-specific data, limiting their scalability and adaptability to new domains or tasks.
-Figure 2: Overview of direct relevance and comparison reasoning prompts for document ranking.
-Pairwise reasoning explains how a document matches a query, while comparison reasoning evaluates the relative relevance between documents.
-Figure 3: Overview of the Reason-to-Rank framework.
+With the explosive growth of online information, modern search engines and recommendation systems heavily rely on reranking to refine initial retrieval results and present users with the most relevant content (Sun et al., 2023; Pradeep et al., 2023b).
+Despite significant progress in neural reranking, particularly with large language models (LLMs) such as GPT-4, contemporary solutions often lack transparency in how they differentiate between highly similar candidate documents.
+Users see a final ranking list, yet remain uncertain about why how specific items are prioritized, which can reduce trust and hinder effective system debugging (Nogueira et al., 2020; Liu et al., 2024).
+Explainable reranking has attracted increasing attention, with the aim of combining ranking decisions with human-readable justifications.
+Recent studies (Yu et al., 2024; Zhuang et al., 2023) show that providing textual explanations can improve user understanding and confidence, and can also help system designers pinpoint potential errors in the ranking pipeline.
+However, most prior approaches focus on a single dimension of explanation, direct relevance reasoning, where each document is described in isolation: “This document is relevant to the query because of…”.
+While useful, such direct rationales do not address the relative importance of documents, a critical aspect when multiple items appear similarly relevant yet differ in subtle ways.
+We argue that comparison reasoning is the missing piece in many existing explainable reranking systems.
+By comparing pairs (or subsets) of documents—explaining, for instance, why Document A is placed above Document B—the model can highlight unique advantages or deficiencies, enabling more fine-grained control over the final ranked list (Pradeep et al., 2021; Sun et al., 2023).
+Without such pairwise or listwise clarity, users and IR practitioners must rely on guesswork when interpreting top results, limiting the overall efficacy of explanation-based retrieval solutions.
+In this paper, we propose Reason-to-Rank (R2R), a novel framework that unifies direct relevance with comparison-based reasoning.
+R2R leverages a powerful teacher LLM (e.g., GPT-4) to produce both a global ranking and two kinds of textual justifications: (1) a direct explanation for each candidate, and (2) a comparative explanation indicating which document outranks another and why.
+The framework is shown in Figure 1.We then employ a knowledge distillation procedure to train a smaller student model to replicate the teacher’s ranking decisions and explanations.
+From an educational perspective, research shows that learners benefit more when they are taught not just answers but the reasoning behind them (Raza, 2020).
+Inspired by this, our approach transfers both the what and the why, helping the student model generalize better and generate its own explanations with improved clarity.
+This setup offers two major advantages: (i) The final deployed system preserves interpretability without incurring the high computational cost of querying a large LLM in production, and (ii) By explicitly encoding comparative rationales, R2R achieves more subtlety distinctions among the top candidates than direct relevance only methods.
+We perform extensive experiments on multiple benchmarks, including MSMARCO (Bajaj et al., 2016), BEIR (Thakur et al., 2021), and BRIGHT (Su et al., 2024), covering both general search queries and reasoning-intensive tasks.
+The results show that R2R outperforms various baselines (e.g., monoT5 (Nogueira et al., 2020), RankVicuna (Pradeep et al., 2023a)) in terms of NDCG metrics, while also providing more interpretable output.
+A comprehensive ablation study further validates the importance of integrating both direct relevance and comparison reasoning, as omitting either dimension leads to decreased performance and reduced clarity.
+In addition, we provide qualitative examples illustrating how comparison explanations help reveal subtle but important differences among retrieved documents.
+In summary, our primary contributions are:
+  • We introduce a novel reranking framework, R2R, that produces both direct relevance and comparison-based explanations, offering a more transparent and fine-grained view of document ranking.
+  • We propose a teacher–student knowledge distillation pipeline, where a large LLM generates enriched multiperspective rationales, subsequently transferred to a smaller, more efficient student model for practical deployment.
+  • We conduct extensive evaluations on general and reasoning-focused datasets, demonstrating that R2R achieves not only competitive or superior reranking performance but also enhanced interpretability through dual-type reasoning.
+Figure 2.
+Overview of direct relevance and comparison reasoning prompts for document ranking.
+Direct relevance reasoning explains how a document matches a query, while comparison reasoning evaluates the relative relevance between documents.
+2. Related Work
+2.1. Neural Reranking and Explanation
+Neural reranking has been extensively studied to refine candidate lists obtained from first-stage retrieval (e.g., BM25 or dense retrievers) (Nogueira et al., 2019, 2020).
+Early efforts often used cross-encoder architectures (e.g., BERT or T5) to capture fine-grained query–document interactions.
+Although these models achieve strong effectiveness, they typically present limited insights regarding the reasoning behind their decisions (Liu et al., 2009).
+Recent work has thus explored generating explanations alongside the ranking (Zhuang et al., 2023; Yu et al., 2024).
+For instance, Explain-then-Rank (Zhuang et al., 2023) adds textual justifications on how a document addresses the query.
+However, these methods tend to focus on direct relevance (i.e., how well a single document matches the query) rather than explicitly modeling comparative relevance among documents, which may be crucial in distinguishing subtle differences at the top of the ranked list (Doshi-Velez and Kim, 2017; Lipton, 2018).
+2.2. Direct vs. Comparison Reasoning
+Classical Learning to Rank (LTR) frameworks (Liu et al., 2009) have shown that pointwise, pairwise, and listwise strategies each possess unique advantages.
+Pairwise approaches like duoBERT (Pradeep et al., 2021) are explicitly designed to compare two documents and decide which one is more relevant, whereas pointwise methods focus on rating each document in isolation.
+More recently, listwise approaches (e.g., RankGPT (Sun et al., 2023) or RankVicuna (Pradeep et al., 2023a)) consider the entire ranked list, often achieving stronger empirical results.
+Nonetheless, even these advanced listwise techniques do not always clearly explain why one document outranks another.
+Our proposed R2R differs by explicitly producing both direct relevance and comparative rationales within a unified framework.
+2.3. Knowledge Distillation in Reranking
+Recent work has addressed the dual goals of ranking effectiveness and transparency.
+On the explanation front, EXS (Singh and Anand, 2019) offers local, model-agnostic rationale extraction for pointwise rankers, whereas RankingSHAP (Heuss et al., 2024) applies Shapley-value analysis across entire ranked lists.
+Both methods clarify why documents are preferred, but do not compress these insights into smaller, more efficient models.
+Meanwhile, methods like One-Shot Labeling (MacAvaney and Soldaini, 2023) and Few-Shot Prompting for Pairwise Ranking (Sinhababu et al., 2024) demonstrate that even sparse or minimal relevance data can significantly boost retrieval.
+Although they refine ranking accuracy via limited supervision, they typically do not target explanation generation or model compression.
+Overall, our approach is complementary: we incorporate both direct and comparative reasoning, and distill these insights into an open-source student model that can both rank and explain.
+Extending Reason-to-Rank to incorporate advanced explanation techniques such as Shapley-based analysis is a promising avenue for future work.
+Figure 3. Basic prompt and return type.
+Figure 4.
+Overview of the Reason-to-Rank framework.
 The teacher model generates direct relevance and comparative reasoning, which is used to train student models capable of reproducing the reranked order and generating explanations.
-Early ranking systems primarily focused on direct relevance methods, where each document’s relevance was judged in isolation Liu et al.
-(2009).
-However, pairwise approaches like duoBERT Pradeep et al.
-(2021) were introduced to compare pairs of documents, determining which document was more relevant.
-While this improved the relative ranking of documents, it still missed the larger picture of the entire ranked list.
-To address this, listwise models such as RankGPT Sun et al.
-(2023) and its distilled versions like RankVicuna Pradeep et al.
-(2023a) considered the entire ranked list during training, allowing for more holistic improvements in ranking accuracy.
-However, even these methods have limitations when it comes to explaining why a particular document ranks higher than another.
-The Learning to Rank framework introduced several foundational approaches in information retrieval, including pointwise, pairwise, and listwise methods, which form the basis of modern neural ranking models Liu et al.
-(2009).
-Direct relevance approaches focus on predicting the relevance score of individual documents in isolation, while pairwise methods aim to indicate the relative ranking between document pairs.
-On the other hand, listwise approaches focus on directly optimizing the entire ranking list, offering better overall accuracy by considering the global structure of the ranking.
-Despite these advances, most early methods lacked clarity that modern applications demand.
-Recent work has also highlighted the importance of natural language explanations in improving clarity and trust in reranking decisions.
-Models such as Explain-then-Rank Zhuang et al.
-(2023) have focused on generating explanations that describe how a document matches a query.
-However, these approaches often only address direct relevance reasoning (i.e., document-query matching) and do not account for comparative rationale (i.e., why one document ranks higher).
-Our work builds on these efforts by incorporating both types of reasoning, enhancing the overall clarity and coherence of the explanations.
-This dual focus provides more precise insights into the ranking process and improves the model’s ability to deliver accurate and logically consistent rankings Doshi-Velez and Kim (2017); Lipton (2018); Mo et al.
-(2024); Yu et al.
-(2024b, a).
-Knowledge distillation has become an essential technique in training smaller, more efficient student models from larger teacher models Hinton et al.
-(2015) to further reduce the computational demands of large reranking models.
-Knowledge distillation has been particularly effective in reranking tasks, as shown by models like RankVicuna Pradeep et al.
-(2023a), which distilled the reranking capabilities of RankGPT into a smaller, more efficient model without significant loss of performance.
-Our approach builds on this by applying distillation to both the reasoning and ranking components, ensuring that our student models not only maintain high performance but also offer explainability in their reranking decisions Qin et al.
-(2023); Dai et al.
-(2022); Xu et al.
-(2024b, a).
-By integrating these advances, our R2R framework improves existing methods by offering a dual reasoning strategy - direct relevance and comparative reasoning - within a computationally efficient model.
-Our approach not only enhances ranking accuracy but also provides clear and easy-to-understand explanations for the ranking decisions, addressing fundamental limitations in performance seen in prior work.
-3 Method
-The overview of the architecture of the R2R model is shown in Figure 3. R2R, a framework that integrates direct relevance and comparative reasoning within a unified reranking approach.
-3.1 Teacher Model with Dual Reasoning
+3. Method
+Initial Retrieval with BM25.
+Following common practice in neural re-ranking (Nogueira et al., 2019; Sun et al., 2023), we first retrieve the top-100 candidate documents per query using BM25.
+This initial retrieval step serves as the candidate pool from which our re-ranking pipeline will operate.
+We choose 100 to balance recall (ensuring that most relevant documents are included) and computational overhead.
+[^1] Our code and data are available here https://github.com/JoyDajunSpaceCraft/Reason-to-Rank.git
+3.1. Teacher Model with Dual Reasoning
 Given a query $q$ and a set of candidate documents $\{d_{1},d_{2},\dots,d_{n}\}$, the teacher model processes these inputs to generate two types of reasoning and rerank order.
@@
 Comparative Reasoning: The model assesses the relative relevance between documents, explaining why one document should be ranked higher based on content.
-3.2 Student Model with Knowledge Distillation
+We also have the basic prompt and the return type in Figure 3.
+Our teacher model (e.g., GPT-4) processes the top-100 candidates and produces two outputs:
+  1. (1)
+    A ranking order (i.e., a global ordering of the 100 documents), which corresponds to how the teacher prioritizes each document’s relevance to the query.
+  2. (2)
+    A textual explanation for each document, covering both direct relevance (i.e., how the document addresses the query) and comparison reasoning (i.e., why one document is more relevant than another).
+While the teacher internally may compute or utilize relevance scores, we unify these into a final ordering to avoid ambiguity.
+In other words, rather than exposing raw numeric scores, the teacher model outputs a definitive rank list (e.g., from most relevant to least relevant) alongside the corresponding rationales.
+3.2. Student Model with Knowledge Distillation
 Deploying large language models (LLMs) for reranking is often impractical due to their high computational cost.
 To address this, we design the student model to replicate the teacher’s performance while being more efficient for real-world deployment.
+We obtain numeric ranking scores for supervision by prompting the teacher to produce a ranked list and converting the document positions into scores via normalized ranks.
 The innovation lies in using knowledge distillation: the student learns from the teacher’s reasoning and ranking outputs, retaining accuracy and interpretability with significantly lower computational demands.
 This approach ensures the student can perform complex reranking tasks efficiently, making it ideal for large-scale use.
-3.2.1 Model Overview
+3.2.1. Model Overview
 Given a query $q$ and a set of candidate documents $\{d_{1},d_{2},\dots,d_{n}\}$, the model processes these inputs to generate both reasoning outputs and ranking scores.
@@
 Ranking Scores: The model simultaneously generates ranking scores $\mathbf{s}_{i}$ for each document, indicating its relevance to the query.
-3.2.2 Training Objective
+3.2.2. Training Objective
 To optimize the performance of the student model, we define three distinct loss functions, each corresponding to different aspects of the model.
@@
 Next, the listwise loss optimizes the overall ranking of the document list.
-It uses Kullback-Leibler (KL) divergence Hershey and Olsen (2007) to measure the difference between the predicted ranking distribution and the target ranking distribution:
+It uses Kullback-Leibler (KL) divergence (Hershey and Olsen, 2007) to measure the difference between the predicted ranking distribution and the target ranking distribution:
 $\mathcal{L}_{\text{listwise}}=D_{\text{KL}}(Q(\mathbf{z})\parallel P(\mathbf{s}))$
 where $P(\mathbf{s})=\text{softmax}(\mathbf{s}_{i})$ is the predicted ranking distribution based on the model’s output, and $Q(\mathbf{z})=\text{softmax}(\mathbf{z}_{i})$ is the target ranking distribution based on true labels.
+We treated the pairwise loss and listwise loss as rank loss.
 Finally, the generation loss optimizes the quality of the textual reasoning generated by the model.
@@
 With this design, the student model efficiently performs ranking tasks and provides transparent reasoning for its decisions, enhancing both transparency and performance.
+The overview of the architecture of the R2R model is shown in Figure 4. R2R, a framework that integrates direct relevance and comparative reasoning within a unified reranking approach.
+Table 1.
+NDCG@5 performance (in percentage) for different models with various reasoning prompts across multiple datasets.
+Bold numbers indicate the best performance within each group of models.
+indicates statistically significant improvement over the base model without reasoning prompts (p $<$ 0.05).
+indicates statistically significant improvement over both the base model and the single reasoning prompt variants (p $<$ 0.05).DL19 and DL20 (Craswell et al., 2020) are test datasets for MSMARCO.
+The MonoT5 has no result for News, and NFCorpus and RankVicuna has no results for News in their paper.
+Models DL19 DL20 BEIR-6 Avg. Covid Touche News NFCorpus Robust04 DBPedia
+BM25 50.6 48.0 41.0 59.5 44.2 39.5 30.8 40.7 31.8
+Rank GPT-4 Model
+Rank GPT-4 (w/o reasoning) 75.6 70.6 54.0 85.5 38.5 57.6 36.6 57.6 47.1
++ direct relevance reasoning 75.3 69.9 52.9 83.3 38.9 55.9 36.1 52.9 47.0
++ comparison reasoning 76.3 71.1 53.4 84.6 38.2 56.5 36.5 56.5 48.2
++ direct relevance & comparison 77.7 73.2 54.4 85.3 38.3 58.4 36.3 58.6 49.5
+Claude Model
+Claude (w/o reasoning) 72.0 68.8 50.8 82.8 36.1 54.4 36.4 50.8 42.5
++ direct relevance reasoning 70.6 68.2 50.3 82.2 35.9 55.4 35.6 52.4 42.5
++ comparison reasoning 71.3 69.8 51.1 83.3 36.4 55.1 36.7 51.5 43.2
++ direct relevance & comparison 72.1 70.0 52.0 84.0 37.3 55.5 36.4 52.7 44.9
+Gemini Model
+Gemini (w/o reasoning) 71.1 68.5 50.1 83.0 35.6 53.7 35.5 50.1 42.9
++ direct relevance reasoning 69.1 67.0 51.1 81.6 34.8 52.9 36.9 51.3 43.9
++ comparison reasoning 70.2 68.2 50.2 82.4 35.9 53.4 35.4 51.1 42.9
++ direct relevance & comparison 71.4 68.6 51.0 83.5 37.0 53.8 36.1 51.8 43.7
+Table 2.
+NDCG@5 performance (in percentage) for student models and baseline comparisons across multiple datasets.
+’w/o Reasoning’ models do not use reasoning prompts and do not generate explanations; they are optimized using the specified ranking loss functions.
+’w/ direct relevance Reasoning’ and ’w/ comparison Reasoning’ models use reasoning prompts and include reasoning loss for generating explanations.
+’’ indicates statistically significant improvement over the baseline DeBERTa model (p $<$ 0.05).
+Bold indicates the best performance among all models for each dataset.
+Entries marked with ’–’ indicate that results are not available for that dataset.
 Models DL19 DL20 BEIR-6 Avg. Covid Touche News NFCorpus Robust04 DBPedia
 Baseline Models
-DeBERTa Sun et al. (2023) 68.5 64.2 47.0 73.4 32.1 50.2 33.7 49.2 45.4
-MonoT5 Nogueira et al. (2020) 74.5 70.4 48.3 80.0 34.1 46.0 35.2
-RankVicuna Pradeep et al. (2023a) 71.1 68.7 48.3 67.1 48.7 38.5 55.7 35.3
-RankZephyr Pradeep et al. (2023b) 72.2 70.5 51.3 85.1 36.5 53.3 38.9 60.7 35.5
-APEER Jin et al. (2024) 74.6 72.3 51.1 83.9 35.3 52.1 33.4 56.0 46.1
+DeBERTa (Sun et al., 2023) 68.5 64.2 47.0 73.4 32.1 50.2 33.7 49.2 45.4
+MonoT5 (Nogueira et al., 2020) 74.5 70.4 48.3 80.0 34.1 46.0 35.2
+RankVicuna (Pradeep et al., 2023a) 71.1 68.7 48.3 67.1 48.7 38.5 55.7 35.3
+RankZephyr (Pradeep et al., 2023b) 72.2 70.5 51.3 85.1 36.5 53.3 38.9 60.7 35.5
+APEER (Jin et al., 2024) 74.6 72.3 51.1 83.9 35.3 52.1 33.4 56.0 46.1
 Our Models (student)
@@
 w/ direct relevance & comparison 75.4 72.4 52.4 84.6 36.2 53.8 36.4 53.5 47.9
-Table 1: NDCG@5 performance (in percentage) for student models and baseline comparisons across multiple datasets.
-’w/o Reasoning’ models do not use reasoning prompts and do not generate explanations; they are optimized using the specified ranking loss functions.
-’w/ direct relevance Reasoning’ and ’w/ comparison Reasoning’ models use reasoning prompts and include reasoning loss for generating explanations.
-’’ indicates statistically significant improvement over the baseline DeBERTa model (p $<$ 0.05).
-Bold indicates the best performance among all models for each dataset.
-Entries marked with ’–’ indicate that results are not available for that dataset.
-4 Experiments
+4. Experiments
 In this section, we evaluate the performance of our Reason-to-Rank (R2R) framework across various information retrieval tasks.
@@
 Finally, we present a comprehensive analysis of the experimental results, including comparisons with existing methods in §4.2.4 and §4.2.1.
-4.1 Datasets and Evaluation Metrics
-In addition to MSMARCO Bajaj et al. (2016) and BEIR Thakur et al. (2021), we utilize the BRIGHT Su et al. (2024) to further evaluate our model’s ability to handle reasoning-intensive retrieval tasks.
+4.1. Datasets and Evaluation Metrics
+In addition to MSMARCO (Bajaj et al., 2016) and BEIR (Thakur et al., 2021), we utilize the BRIGHT (Su et al., 2024) to further evaluate our model’s ability to handle reasoning-intensive retrieval tasks.
 Evaluation Metrics
@@
 BLEU, and ROUGE-L: Used for evaluating the quality of generated reasons in reasoning tasks, assessing both lexical overlap and logical consistency.
-4.2 Models
-4.2.1 Teacher models
-We utilize large language models (LLMs) as teacher models to generate reasoning annotations: GPT-4 Achiam et al.
-(2023), Claude The (claude-3-5-sonnet-20240620) and Gemini Team et al.
-(2023) (gemini-1.5-flash).
+4.2. Models
+4.2.1. Teacher models
+We utilize large language models (LLMs) as teacher models to generate reasoning annotations: GPT-4 (Achiam et al., 2023), Claude (The, [n.
+d.]) (claude-3-5-sonnet-20240620) and Gemini (Team et al., 2023) (gemini-1.5-flash).
 These advanced LLMs known for their strong reasoning and language-understanding capabilities.
@@
 Given GPT-4’s overall solid performance and the boost from reasoning prompts, we select it as the teacher model to distill its enhanced reasoning abilities into our student models.
-4.2.2 Student models
+4.2.2. Student models
 We implement our models using LoRA (Low-Rank Adaptation) on 32GB V100 GPUs to fine-tune the student models on large datasets efficiently.
@@
 Impact of Reasoning Strategies
-we also evaluated different reasoning strategies in the R2R student models, specifically analyzing the "w/o Reasoning" (without reasoning), "w/ direct relevance reasoning" (with direct relevance reasoning), "w/ comparison reasoning" (with comparison reasoning), and "w/ direct relevance & comparison" (with both direct relevance and comparison reasoning) variants.
-As shown in Table  1, models without reasoning (using pairwise or listwise loss) perform at a baseline level, with NDCG@5 scores of 73.2% and 73.8 on DL19 and 70.8% and 71.0% on DL20.
+we also evaluated different reasoning strategies in the R2R student models, specifically analyzing the ”w/o Reasoning” (without reasoning), ”w/ direct relevance reasoning” (with direct relevance reasoning), ”w/ comparison reasoning” (with comparison reasoning), and ”w/ direct relevance & comparison” (with both direct relevance and comparison reasoning) variants.
+As shown in Table  2, models without reasoning (using pairwise or listwise loss) perform at a baseline level, with NDCG@5 scores of 73.2% and 73.8% on DL19 and 70.8% and 71.0% on DL20.
 Adding reasoning significantly boosts performance.
@@
 Interestingly, while reasoning improves performance in most cases, there are instances where models without reasoning (e.g., listwise loss) outperform those with reasoning.
-4.2.3 Comparison with Baselines
+4.2.3. Comparison with Baselines
 Our student models, trained with direct relevance and comparison reasoning, achieve competitive performance compared to the previous state-of-the-art student models.
-As shown in Table  1,the improvements over the baseline models are statistically significant, with p-values less than 0.05.
+As shown in Table  2,the improvements over the baseline models are statistically significant, with p-values less than 0.05.
 BM25: A traditional term-based retrieval model using TF-IDF, known for its simplicity and efficiency in keyword-based retrieval.
-RankGPT DeBERTa Sun et al. (2023): A transformer model using disentangled attention to better capture long-range dependencies, improving reranking by understanding query-document context.
-MonoT5 Nogueira et al. (2020): A T5-based model treating reranking as a sequence generation task, leveraging its encoder-decoder architecture for better relevance scoring.
-RankVicuna, RankZephyr and APEER Pradeep et al.
-(2023a, b); Jin et al.
-(2024): RankVicuna is a distilled version of RankGPT, designed for zero-shot listwise reranking with fewer parameters, balancing efficiency and performance.
+RankGPT DeBERTa (Sun et al., 2023): A transformer model using disentangled attention to better capture long-range dependencies, improving reranking by understanding query-document context.
+MonoT5 (Nogueira et al., 2020): A T5-based model treating reranking as a sequence generation task, leveraging its encoder-decoder architecture for better relevance scoring.
+RankVicuna, RankZephyr and APEER (Pradeep et al., 2023a, b; Jin et al., 2024): RankVicuna is a distilled version of RankGPT, designed for zero-shot listwise reranking with fewer parameters, balancing efficiency and performance.
 RankZephyr a state-of-the-art zero-shot listwise reranking model that adapts RankGPT prompts for robust performance with minimal task-specific training.
 APEER share the same structure with RankZephyr but have an automated prompt generation for large language models.
-4.2.4 R2R Student Model vs. Baselines
+4.2.4. R2R Student Model vs. Baselines
 Our R2R student model indicates strong performance across various datasets, often outperforming the baseline models.
@@
 Similarly, on the Robust04 dataset, RankZephyr attains a higher NDCG@5 of 60.7% compared to our model’s 53.5%.
-4.2.5 Performance on Reasoning-Intensive Tasks (BRIGHT Dataset)
-The BRIGHT dataset Su et al.
-(2024) is specifically designed to evaluate models on reasoning-intensive retrieval tasks, simulating real-world scenarios where complex reasoning is required to determine document relevance.
-We report the Inst-LSu et al.
-(2024, 2022) model as a baseline because it is fine-tuned with task-specific instructions, making it particularly well-suited for tasks requiring deeper reasoning.
+For Touche, it is shown that it is a argument retrieval or debate scenario dataset, and BM25 tends to perform surprisingly well in such “keyword-matching” strong correlation scenarios; furthermore, some models (e.g., RankVicuna) may be better adapted in argument mining Furthermore, some models (e.g., RankVicuna) may be better adapted for argument mining.
+For Robust04, this means that it covers a wider range of topics and outdated documents, where RankZephyr’s multi-iteration listwise training is more advantageous.
+4.2.5. Performance on Reasoning-Intensive Tasks (BRIGHT Dataset)
+The BRIGHT dataset (Su et al., 2024) is specifically designed to evaluate models on reasoning-intensive retrieval tasks, simulating real-world scenarios where complex reasoning is required to determine document relevance.
+We report the Inst-L(Su et al., 2024, 2022) model as a baseline because it is fine-tuned with task-specific instructions, making it particularly well-suited for tasks requiring deeper reasoning.
 Table 3 shows the NDCG@5 results on the BRIGHT dataset.
@@
 However, we observe that in certain domains, such as Pony and AoPS, the performance of our student and teacher models is relatively low.
-Domain BM25 Inst-L Su et al. (2024) R2R GPT-4 Teacher
+Table 3. NDCG@5 results on the BRIGHT dataset.
+Domain BM25 Inst-L(Su et al., 2024) R2R GPT-4 Teacher
 Bio. 17.2 21.0 24.3 28.0
@@
 Average 13.1 17.5 19.6 23.2
-Table 2: NDCG@5 results on the BRIGHT dataset.
-5 Ablation Study
-5.1 Quality of Generated Reasons
+5. Ablation Study
+5.1. Quality of Generated Reasons
 We assess the quality of the generated reasons using BLEU and ROUGE-L as presented in Table 4 compared with the teacher model’s reason.
 The baseline model here represents the model trained without the rank loss function and has the same input as ours.
-On the DL19 dataset, the student model achieves a BLEU score of 21.4, compared to the baseline’s 18.0, and a ROUGE-L score of 36.8 versus 32.5.
-Similar improvements are observed on DL20, where the student model attains BLEU and ROUGE-L scores of 24.0 and 38.2, respectively, surpassing the baseline’s 19.5 and 34.0.
+On the DL19 dataset, the student model achieves a BLEU score of 21.4%, compared to the baseline’s 18.0%, and a ROUGE-L score of 36.8% versus 32.5%.
+Similar improvements are observed on DL20, where the student model attains BLEU and ROUGE-L scores of 24.0% and 38.2%, respectively, surpassing the baseline’s 19.5% and 34.0%.
+Table 4.
+Evaluation of reason generation using BLEU and ROUGE-L scores for student models trained with and without the ranking loss function.
+The ’Student Model w/o Rank Loss’ is trained without the rank loss function and has the same input as the ’Student Model w/ Rank Loss’.
 Dataset Student Model BLEU ROUGE-L
@@
 w/ Rank Loss 22.7 39.3
-Table 3: Evaluation of reason generation using BLEU and ROUGE-L scores for student models trained with and without the ranking loss function.
-The ’Student Model w/o Rank Loss’ is trained without the rank loss function and has the same input as the ’Student Model w/ Rank Loss’.
-5.2 Impact of Reasoning Strategies
-In Table 4, by removing comparison reasoning results in a performance drop on DL19 from 75.3 to 74.1 (NDCG@5).
-Similarly, removing direct relevance reasoning causes a performance decrease on DL20 from 72.3 to 72.2 (NDCG@5).
+5.2. Impact of Reasoning Strategies
+In Table 5, by removing comparison reasoning results in a performance drop on DL19 from 75.3% to 74.1% (NDCG@5).
+Similarly, removing direct relevance reasoning causes a performance decrease on DL20 from 72.3% to 72.2% (NDCG@5).
 Both reasoning strategies contribute to performance improvements.
+Table 5. Ablation study: Impact of direct relevance and comparison reasoning on NDCG@5/@10 for DL19 and DL20.
 Model DL19 NDCG@5/@10 DL20 NDCG@5/@10
@@
 w/o reasoning 73.2 / 72.1 70.8 / 69.8
-Table 4: Ablation study: Impact of direct relevance and comparison reasoning on NDCG@5/@10 for DL19 and DL20.
-5.3 Effects of Training Data Size
-Figure 4 i shows the performance of small models trained on 100, 1,000, and 2,000 examples.
+5.3. Effects of Training Data Size
+Figure 5 shows the performance of small models trained on 100, 1,000, and 2,000 examples.
 Contrary to expectations, increasing data size from 1,000 to 2,000 does not consistently improve performance and sometimes slightly reduces it.
-We find similar trends across various small models, including Mistral and LLaMA variants, highlighting that adding more data does not always guarantee better performance.
-Figure 4: Student model performance for different training data sizes on DL19 and DL20.
-Models DL19 DL20 BEIR-6 Avg. Covid Touche News NFCorpus Robust04 DBPedia
-BM25 50.6 48.0 41.0 59.5 44.2 39.5 30.8 40.7 31.8
-Rank GPT-4 Model
-Rank GPT-4 (w/o reasoning) 75.6 70.6 54.0 85.5 38.5 57.6 36.6 57.6 47.1
-+ pairwise reasoning 75.3 69.9 52.9 83.3 38.9 55.9 36.1 52.9 47.0
-+ comparison reasoning 76.3 71.1 53.4 84.6 38.2 56.5 36.5 56.5 48.2
-+ direct relevance & comparison 77.7 73.2 54.4 85.3 38.3 58.4 36.3 58.6 49.5
-Claude Model
-Claude (w/o reasoning) 72.0 68.8 50.8 82.8 36.1 54.4 36.4 50.8 42.5
-+ pairwise reasoning 70.6 68.2 50.3 82.2 35.9 55.4 35.6 52.4 42.5
-+ comparison reasoning 71.3 69.8 51.1 83.3 36.4 55.1 36.7 51.5 43.2
-+ direct relevance & comparison 72.1 70.0 52.0 84.0 37.3 55.5 36.4 52.7 44.9
-Gemini Model
-Gemini (w/o reasoning) 71.1 68.5 50.1 83.0 35.6 53.7 35.5 50.1 42.9
-+ pairwise reasoning 69.1 67.0 51.1 81.6 34.8 52.9 36.9 51.3 43.9
-+ comparison reasoning 70.2 68.2 50.2 82.4 35.9 53.4 35.4 51.1 42.9
-+ direct relevance & comparison 71.4 68.6 51.0 83.5 37.0 53.8 36.1 51.8 43.7
-Table 5: NDCG@5 performance (in percentage) for different models with various reasoning prompts across multiple datasets.
-Bold numbers indicate the best performance within each group of models.
-indicates statistically significant improvement over the base model without reasoning prompts (p $<$ 0.05).
-indicates statistically significant improvement over both the base model and the single reasoning prompt variants (p $<$ 0.05).
-6 Discussion
-Our experiments show that the R2R framework effectively enhances document reranking by incorporating direct relevance and comparative reasoning.
-The student models achieve competitive results while significantly reducing computational resources compared to large teacher models.
-The statistically significant improvements confirm that reasoning strategies effectively enhance model performance.
-An important aspect of R2R is its ability to not only rank documents but also generate coherent and relevant reasoning for its rankings.
-This dual capability suggests that the model could be applied to other tasks requiring both ranking and reasoning generation.
-While our current work centers on document reranking in information retrieval, the integration of reasoning with ranking decisions holds promise for broader applications.
-For instance, in recommendation systems, the model could rank products or services while providing explanations for its choices, enhancing user trust.
-Similarly, in question answering or summarization tasks, generating reasoning alongside answers or summaries could improve understandability Jiang et al.
-(2024).
-However, future work is needed to explore the model’s performance in these varied contexts and verify its generalizability beyond document reranking McElfresh et al.
-(2022); Rahmani et al.
-(2022).
-7 Conclusion
-In this paper, we introduced ReasoningRank, a novel approach to document reranking that incorporates direct relevance and comparative reasoning.
-Our method leverages the power of large language models to generate interpretable reasons, which are then distilled into smaller, more efficient student models.
-Our experiments show the effectiveness of this approach, achieving high performance across multiple datasets while maintaining interpretability.
-Limitations
-While our approach shows strong performance, it has several limitations.
-First, the reliance on large language models for generating reasoning limits the applicability of our method in resource-constrained environments.
-Second, the quality of the generated reasons heavily depends on the quality of the prompts, which requires careful crafting.
-Future research should address these limitations by exploring more efficient models and automated prompt-generation techniques.
-Ethics Statement
-This work enhances the transparency and accessibility of reasoning capabilities in AI by distilling knowledge from LLMs to smaller models.
-While this improves efficiency, we acknowledge the ethical concerns of bias in LLMs and the need for careful mitigation in real-world applications.
-Appendix A Student Training Parameters
+Similar trends across various small models, including Mistral and LLaMA variants, highlight that adding more data does not always guarantee better performance.
+Figure 5. Student model performance for different training data sizes on DL19 and DL20.
+Table 6.
+Impact of different $(\alpha,\beta,\gamma)$ combinations on model performance.
+We report the NDCG (averaged at a certain cutoff), BLEU, and ROUGE-L scores under different weighting strategies.
+The best performance is in bold.
+Hyperparameters Performance
+$\alpha$ $\beta$ $\gamma$ NDCG BLEU ROUGE-L
+0.2 0.2 0.6 0.78 0.74 0.72
+0.3 0.3 0.5 0.82 0.76 0.75
+0.4 0.4 0.2 0.86 0.79 0.78
+0.5 0.5 0.1 0.84 0.77 0.76
+0.6 0.6 0.05 0.80 0.73 0.74
+5.4. Hyperparameter Tuning.
+We conducted a small-scale grid search over $\alpha$, $\beta$,$\gamma$ on a held-out validation set.
+Table 6 shows the results for different weighting strategies.
+We find that ($\alpha$=0.4,$\beta$=0.4,$\gamma$=0.2) yields the best trade-off among NDCG, BLEU, and ROUGE-L, and we thus adopt these settings for all subsequent experiments.
+5.5. Case Study
+In addition to ranking improvements, the generated explanations offer interpretability benefits.
+For example, in several case studies, see Table 7, we observed that the comparison reasoning often highlights subtle differences between top-ranked documents, which can assist developers in diagnosing ranking errors.
+Although our current evaluation of explanations relies on automated metrics (BLEU, ROUGE), a small-scale human evaluation is planned in future work to quantitatively assess readability, consistency, and practical utility of these explanations.
+Table 7.
+An illustrative case study showing how direct and comparison reasoning enhance interpretability for a medical-related query.
+In this example, Document A is ultimately deemed more relevant due to its detailed coverage of insulin’s regulatory mechanism, whereas Document B offers a more general discussion of hormones.
+Query: ”How does insulin regulate blood sugar?”
+Document A (snippet)
+“Insulin is produced by the pancreas and helps cells take in glucose from the bloodstream. By binding to insulin receptors, it initiates a signaling cascade that lowers blood glucose levels…”
+Direct Reasoning:
+“This document directly addresses how insulin facilitates cellular glucose uptake, which aligns precisely with the question on regulating blood sugar.”
+Comparison Reasoning vs. Document B:
+“Compared to Document B, Document A provides more explicit details on insulin’s binding mechanism and the biochemical pathway, making it more relevant for understanding exactly how insulin regulates glucose.”
+Document B (snippet)
+“Hormones like insulin and glucagon maintain metabolic balance. Insulin generally lowers blood sugar levels, while glucagon raises them, ensuring homeostasis…”
+Direct Reasoning:
+“This document briefly mentions insulin’s role in lowering blood sugar, but primarily focuses on the broader hormone interplay.”
+Comparison Reasoning vs. Document A:
+“Although Document B is relevant, it emphasizes multiple hormones and offers less specific detail on the mechanism of insulin.
+Document A is therefore more closely aligned with how insulin regulates blood sugar.”
+6. Discussion and Conclusion
+Our experimental results demonstrate that Reason-to-Rank (R2R) effectively enhances document reranking by integrating both direct relevance and comparative reasoning.
+Compared to relying solely on a large teacher model (e.g., GPT-4), our distilled student models offer competitive performance while drastically reducing computational cost, making the framework suitable for real-world, high-throughput settings.
+The statistically significant gains confirm that incorporating explicit reasoning structures—both pointwise and pairwise—meaningfully contribute to improved ranking accuracy.
+A natural concern is whether the observed performance improvements stem merely from generating longer output text rather than the intrinsic value of direct and comparison reasoning.
+We argue that the benefit of our approach lies in the structured content of the explanations: by explicitly contrasting documents (comparison reasoning) and detailing query-document alignment (direct reasoning), the model captures fine-grained distinctions that standard ranking scores might overlook.
+Although we have not conducted a controlled experiment to equalize text length, preliminary analyses indicate that even when constraining the output to a fixed token budget, models incorporating both reasoning types consistently outperform those without.
+Future work will include a more rigorous ablation study and a human evaluation to further assess the usefulness of the generated rationales in aiding system diagnosis and debugging.
+Notably, the proposed student models not only output document rankings but also produce coherent and relevant justifications for each document’s position.
+This dual capacity to rank and explain suggests broader applicability beyond traditional ad-hoc retrieval.
+For example, in recommendation systems, R2R could prioritize items while elucidating the rationale (e.g., product features or user context), thereby increasing user trust and satisfaction.
+Similarly, in question answering or summarization tasks, generating reasoning alongside answers may boost transparency and user acceptance (Jiang et al., 2024).
+However, systematically evaluating these broader domains remains as future work to verify the generalizability of our approach (McElfresh et al., 2022; Rahmani et al., 2022).
+Overall, the synergy between direct relevance and comparative reasoning leads to both practical efficiency and rich interpretability.
+Our findings point to an emerging trend in IR and NLP systems where explicit reasoning improves performance metrics and contributes to more explainable and user-friendly outcomes.
+In this paper, we introduced Reason-to-Rank (R2R), a novel framework that unifies direct relevance and comparison reasoning for document reranking.
+We leveraged large language models (LLMs) as teacher models to generate high-quality explanations and rankings, then distilled both the ranking expertise and textual justifications into smaller student models.
+Experimental results across multiple benchmarks illustrate that R2R not only achieves strong reranking effectiveness but also delivers interpretable rationales for why one document is ranked above another.
+Our study underscores the value of distilling explicit reasoning strategies from large LLMs and highlights the benefits of merging pointwise and pairwise perspectives.
+7. Limitations and Future Work
+Despite the encouraging results, there are several limitations to consider: Dependence on large LLMs for teacher data.
+Our approach relies on powerful teacher models (e.g., GPT-4) to generate high-quality explanations.
+Obtaining robust supervision signals could be challenging in resource-constrained environments or specialized domains where such models are not readily available.
+Also, using GPT-4 can be expensive (approx.\$0.2 per query with 4,000+ tokens).
+While knowledge distillation substantially reduces cost at inference time, the distilled models are still large and thus not trivial for real-time production in specific scenarios.
+Prompt quality and model biases.
+The effectiveness of the distilled reasoning hinges on carefully designed prompts.
+If prompts fail to capture key aspects of the query or document, the generated explanations might be shallow or incorrect.
+Additionally, teacher models can introduce biases, which might propagate into student models during distillation.
+Broader applicability beyond reranking.
+While we have discussed possible extensions to recommendation, QA, or summarization tasks, the actual performance and adaptation strategy in these domains remain largely unexplored.
+Thorough empirical studies are required to confirm that the proposed dual reasoning approach generalizes effectively.
+Lack of Human Evaluation for Explanations We only measured textual explanations with BLEU/ROUGE against the teacher outputs, which themselves are automatically generated.
+This does not guarantee that the explanations are genuinely helpful or correct from a human perspective.
+Generalization beyond Document Ranking.
+Though we focus on ad-hoc retrieval, the idea of distilling direct and comparative reasoning could benefit other tasks like question answering or recommender systems.
+More thorough domain-specific experiments are necessary to confirm the general applicability of our approach.Confounding Factors of Explanation Length.While our ablation controls for token length, further studies are needed to rigorously isolate how much of the ranking improvement stems from the reasoning structure vs.
+extended context.
+Understanding how best to prompt for concise but effective rationales remains an open question.
+In future work, we plan to investigate more efficient teacher–student paradigms (e.g., smaller teacher models, automated prompt generation) and conduct domain-specific evaluations to further validate the versatility of R2R.
+We believe these efforts can refine the balance between interpretability and computational cost, extending the framework’s applicability across diverse IR and NLP scenarios.
+8. Appendix
+8.1. Student Training Parameters
+Table 8. Training configuration hyperparameters.
 Hyperparameter Value
@@
 freeze_layers False
-Table 6: Training configuration hyperparameters.
-Appendix B Teacher Model Parameters
-Hyperparameter Value
-Model GPT
-Temperature 1.0
-Top p 0.9
-Table 7: Teacher model hyperparameter settings.
-Appendix C NDCG@5 for Different Models
-Table 8: NDCG@5 for different models and training data sizes on various datasets.
-Model Train Size DL 19 DL 20
-Mistral v0.1 100 69.69 76.33
-1000 71.03 77.06
-2000 70.63 73.99
-Mistral v0.3 100 69.26 79.55
-1000 71.56 79.14
-2000 69.86 76.96
-LLaMA 3.1 8B 100 69.63 75.85
-1000 72.87 77.03
-2000 73.37 77.60
-LLaMA 3 8B 100 69.57 77.15
-1000 73.76 75.35
-2000 71.75 75.55
-LLaMA 2 7B 100 71.50 79.89
-1000 72.19 72.95
-2000 72.18 75.36
-Appendix D Model Behavior Analysis
-In the reasoning generation tasks, the model’s performance varied across direct relevance reasoning, comparison reasoning, and their combination.
-Below, we provide an analysis of unexpected behavior based on 2,000 queries.
-Repetition.
-The repetition metric measures the occurrence of duplicate passage identifiers. Across the tasks, we find:
-  • Pairwise reasoning: 63 duplicate passage identifiers (3%).
-  • Comparison reasoning: 36 duplicate passage identifiers (1.5%).
-  • Combined reasoning: 47 duplicate passage identifiers (2%).
-Missing Documents.
-Missing documents were most frequent in direct relevance reasoning:
-  • Pairwise reasoning: 642 missing documents (32% of queries).
-  • Comparison reasoning: 227 missing documents (11% of queries).
-  • Combined reasoning: 328 missing documents (16% of queries).
-Appendix E API Cost Analysis
-Utilizing large language models (LLMs) like GPT-4 as teacher models involves significant computational and monetary costs. In this section, we analyze these costs based on our experiments.
-We calculated the average number of tokens, the number of API requests, and the estimated cost per query for different operations involving the teacher model.
-The costs are based on the pricing provided by the API service at the time of our experiments 9.
+8.2. API Cost Analysis
+We calculated the average number of tokens, the number of API requests, and the estimated cost per query for different operations involving the teacher model, see in Table 9.
+The costs are based on the pricing provided by the API service at the time of our experiments.
+Table 9. Average token usage and cost per query using GPT-4 for different reasoning tasks.
 API Instruction Tokens Cost (\$USD)
@@
 GPT-4 Direct relevance & Comparison 4650 0.194
-Table 9: Average token usage and cost per query using GPT-4 for different reasoning tasks.
-Appendix F Prompts for Reasoning
-Table 10: Basic prompt, direct relevance reasoning prompt, and comparison reasoning prompt.
-Basic Prompt:
-I will provide you with num passages, each indicated by a number identifier [].
-For each passage, briefly generate your reasoning process as follows:
-1. Judge whether the passage is not applicable (not very common), where the query does not meet the
-premise of the passage.
-2. Check if the query contains direct evidence. If so, judge whether the query meets or does not meet
-the passage.
-3. If there is no direct evidence, try to infer from existing evidence and answer one question:
-If the passage is ranked in this order, is it possible that a good passage will miss such information?
-If impossible, then you can assume that the passage should not be ranked in that order.
-Otherwise, it should be ranked in that order.
-Direct relevance Reasoning Prompt:
-Then, read every passage one by one and find the sentence where the method is the direct answer
-for the query and output the sentence. Then, rank the passages based on their relevance to the query.
-Give highest priority to passages that provide a clear and direct definition or explanation of the
-keywords related to the query.
-Consider both the detailed information and any relevant background context provided in each passage.
-Provide clear and concise reasons for the ranking, highlighting the specific parts of the passages that
-influenced your decision.
-Make sure to ignore any irrelevant information and focus on content directly related to the query.
-In addition to direct reasons for ranking each passage, also consider listwise reasons.
-Comparison Reasoning Prompt:
-A listwise reason involves comparing passages with each other to determine their relative importance.
-Specifically, compare passages to identify which ones provide more comprehensive, relevant, and accurate
-information in relation to the query.
-When providing listwise reasons, mention specific comparative insights, such as why one passage might be
-more relevant than another based on the overall context and detail provided.
-Return Type:
-Search Query: query. Rank the num passages above based on their relevance to the search query and
-the extracted keywords. The passages should be listed in descending order using identifiers.
-The most relevant passages should be listed first. The output should be in JSON format.
-For each passage, provide the extracted keywords and detailed reasons for its ranking.
-Ensure to separate direct reasons and listwise reasons clearly.
-Mention specific parts of the passage that influenced your decision.
-Ensure the reasons are clear, concise, and directly related to the query,
-balancing direct definitions or explanations and relevant background information.
-Only output the JSON structured format.